### v1

# Accelerating Quantum Eigensolver Algorithms With Machine Learning

Abstract
In this paper, we explore accelerating Hamiltonian ground state energy calculation on NISQ devices. We suggest using search-based methods together with machine learning to accelerate quantum algorithms, exemplified in the Quantum Eigensolver use case. We trained two small models on classically mined data from systems with up to 16 qubits, using XGBoost’s Python regressor. We evaluated our preliminary approach on 20-, 24- and 28-qubit systems by optimising the Eigensolver’s hyperparameters. These models predict hyperparameter values, leading to a 0.13%-0.15% reduction in error when tested on 28-qubit systems. However, due to inconclusive results with 20- and 24-qubit systems, we suggest further examination of the training data based on Hamiltonian characteristics. In future work, we plan to train machine learning models to optimise other aspects or subroutines of quantum algorithm execution beyond its hyperparameters.
1 Introduction
Modern-day Noisy Intermediate-Scale Quantum (NISQ) devices are the current state-of-the-art in Quantum Computing (QC) characterised as noisy with an intermediate scale in terms of the number of qubits they have [43, 10, 1]. These devices have not yet evolved to support fault-tolerant calculations with a large enough number of qubits to achieve quantum advantage. However, these already enable quantum researchers and engineers to develop, optimise and test various quantum algorithms. In this interim time of intensive hardware development, robust and innovative quantum algorithms [34] provide a way to tackle the limitations of the current quantum processors. One of the most promising applications of QC is in the field of quantum chemistry for molecular simulations, structure design, drug design and more [37, 33, 35]. A prominent example of a class of quantum algorithms that is rapidly evolving in the field of quantum chemistry is the Variational Quantum Eigensolver (VQE) [51], which aims at estimating the ground state energy of Hamiltonians. Different types of VQE algorithms are being developed and tailored for various systems.
Whilst taking part in the Quantum Algorithm Grand Challenge [46] organised by QunaSys [^1] QunaSys is a technology company aiming to achieve QC’s full potential through algorithm development into product engineering [44], we tackle an instance of quantum algorithms for eigensolving, designed to estimate the eigenvalues and eigenvectors of a given Hamiltonian to find its ground state energy. Our approach focuses on applying machine learning algorithms to explore and optimise quantum algorithms to achieve potential advancements in quantum eigensolving and improve the performance and accuracy of quantum algorithms for larger system setups that are currently too complex to tackle. Our first attempt at the problem was to find a methodology to select the best hyperparameters for the ADAPT-QSCI quantum algorithm [32] that minimise the runtime and the final result for the ground state energy through using machine learning techniques trained on smaller systems [13]. This method is not specifically tailored to ADAPT-QSCI and can be applied to any Quantum Eigensolver-based algorithm. Next, we applied the Quantum Complex Exponential Least Squares (QCELS) algorithm [17] to the problem. We investigated the suitability and limitations of the QCELS algorithm when applied to large systems [14]. In the evaluation, we have assessed their performance and potential benefits for solving 16 quantum systems of 20-, 24- and 28-qubits, showing limited minor improvement for 20- and 28-qubit systems.
The rest of the paper is structured as follows. Section 2 lists related work in the field and Section 3 focuses on the background of the quantum computing algorithms that we have investigated, software engineering optimisation methods and machine learning, followed by a deeper dive into the methodologies of quantum eigensolver algorithms in Section 4, and machine learning applications to accelerate quantum algorithms in Section 5. In Section 6, we present the experiment’s methodology and our results. Our conclusions can be found in Section 7, including notes containing references to the source code used in the paper and the relevant data and acknowledgements.
A note on the 2024 Quantum Algorithm Grand Challenge (QAGC2024).
The challenge focused on improving current quantum algorithms for the problem of finding the ground state energy of a one-dimensional orbital rotated Fermi-Hubbard model Hamiltonian using a given 28-qubit system in the most optimal way [46]. The aim of the challenge is to design an algorithm, given specific constraints (i.e. the number of shots of $10^{7}$, timeout of $6\times 10^{5}$ seconds and limited input data), that can output the closest answer to the actual ground state energy. Previous year’s challenge [45] had the same aim but was based on a much smaller 8-qubit system. The winning algorithm was the company’s proposition later published as Quantum-Selected Configuration Interaction (QSCI) [32].
2 Related Work
Quantum Computing.
Quantum Computing (QC) was conceptualised initially to simulate quantum mechanics using computers “built of quantum mechanical elements which obey [the] quantum mechanical law” [22]. Later, it was found that QC could have several potential applications and offer significant speed-up over classical computing [5, 4, 16, 6, 27, 11]. In 1994, Shor’s proposal of a polynomial-time algorithm for prime factorization and discrete logarithms on a quantum computer raised enormous interest due to its potential threat to modern RSA cryptosystems [49]. Soon after, Grover introduced a fast database search on quantum computers that promised quadratic speed-up over the best classical algorithm [27]. The resulting potential speed-up is often referred to as “quantum supremacy” [3]. Several studies have applied software engineering techniques to optimise quantum computing [24, 25], while testing and debugging approaches have been found to be beneficial in quantum software development [39, 59].
NISQ.
Demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where quantum devices would solve real-life calculations. Although quantum supremacy seems difficult to achieve soon, NISQ algorithms are a prominent example that hybrid systems combining small quantum circuits with classical computations could present some computational advantages, i.e., a quantum advantage [43]. Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era [43]. Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop [51].
Machine Learning.
Machine Learning (ML) algorithms are increasingly used to improve and automate software engineering tasks [55, 58, 28, 48, 19], especially after the advent of Large Language Models (LLM), with common applications in software engineering, including optimisation, code generation, bug detection and automated testing [20, 53, 12, 9, 15, 30]. Furthermore, connections between Machine Learning and QC have been broadly explored, both to optimise Machine Learning with QC and to optimise QC with Machine Learning [21, 41, 54]. As we aim to optimise the quantum algorithm hyper-parameters classically, we discuss how we apply machine learning to enhance the software engineering aspects of our approach in Section 5.
3 Background
Our approach combines machine learning, search-based software engineering, and quantum physics to accelerate quantum algorithms, specifically eigensolvers. We provide background on each area before discussing the applications of machine learning in quantum algorithms. In Section 4, we discuss the implementation of the eigensolvers used in this research.
3.1 Variational Quantum Algorithms
Variational Quantum Algorithms (VQA) are among the most promising examples of NISQ algorithms [51]. The main goal of a VQA is to find the optimal parameters for a parameterized quantum circuit, leading to a solution for a given computational problem. We provide an overview of the structure and functioning of a VQA.
  1. 1.
    Problem definition. The first step involves defining a computational problem that can benefit from quantum processing and translating it into an objective or cost function. In most applications, it consists of a Hamiltonian construction and representation of a quantum system.
  2. 2.
    Optimisation process. A parameterized quantum circuit, known as the variational ansatz, is designed. This circuit contains gates with adjustable parameters, denoted as $\theta$. A quantum state |$\ket{\psi(\theta)}$ is measured, and the outcomes are used to compute the expectation value of the cost function $\bra{\psi(\theta)}H\ket{\psi(\theta)}$. H is the Hamiltonian of the quantum system, and the expectation value of this Hamiltonian is the cost function.
  3. 3.
    Convergence check and output The optimisation process continues until a convergence criterion is met, indicating that further iterations are unlikely to significantly improve the solution. This convergence check ensures that the algorithm has reached a stable and potentially optimal solution. The final set of optimised parameters $\theta_{opt}$ represents the solution to the quantum problem. This solution can be used for further analysis or as the output of the VQA.
An in-depth explanation and a detailed figure can be found in the work of Bharti et al. [10].
VQE and ADAPT-VQE.
Variational Quantum Eigensolvers (VQE) were introduced in 2014 by Peruzzo et al. [42]. It is the most popular VQA, and aims to approximate the ground state energy of a quantum system iteratively. Although considered among the most promising NISQ algorithms, VQEs present several limitations. VQE optimisations have been widely investigated, from classical optimisers, to measurement strategies and Ansatz structure [51]. One popular algorithm derived from VQEs is the ADAPT-VQE [26]. The main distinction is the restructuring of the Ansatz at each iteration. A pool of operators is defined from which operators are selected to update the ansatz during the optimisation process. ADAPT-VQEs are found to have more precise results in some applications, at a cost of a higher computational load.
3.2 Use of AI with GI and Search-based Optimisation
In this work, we integrate ML with search-based methods. We first use predictions on smaller systems, utilising the cost function values of a Hamiltonian of smaller systems and randomly generated hyperparameters of the quantum algorithm to train a gradient boosting model to predict the cost function value. We then use this model to explore and optimise hyperparameters for a larger Hamiltonian system, which is costly to solve classically, via a crossover operator.
Search-based Optimisation and Genetic Improvement (GI).
Search-based methods [29] have also been used for optimisation tasks to find the best possible subset of requirements. The optimisation, in the case of the quantum algorithm’s hyperparameters, aims to save precious resources like the number of shots but mainly to achieve a closer prediction to the system’s ground state energy [^2] The lowest energy level, which for eigensolvers is the lowest eigenvalue that is the value associated with the lowest eigenvector of the problem, see Section 4., influenced by the selection of optimal hyperparameters (i.e. a minimum of the cost function approximating the lowest energy level, Subsection 3.1).
These methods aim to find the near-optimal solution iteratively, ending when a stopping condition is met (e.g. reaching a maximum number of iterations). Starting with a set of candidate solutions, they are evaluated using a fitness function. GI then diversifies solutions over iterations through mutation operations like bit flip (on a single solution) or crossover (on multiple solutions) to potentially generate better-performing offspring. The method can occasionally minimise the population or inject noise to avoid converging to local minima (or maxima, depending on the problem). Using ML to estimate the fitness function is common when direct evaluation or computation is infeasible. We utilise this idea for larger Hamiltonians in Section 5.
In the context of this work, we define a solution as an assignment of hyperparameters (their values based on some pre-defined constraints, e.g. sampling_shots are between $100$ to $10^{6}$), with the best solution being a global minimum. Consequently, our fitness function is related to the lowest energy level for a specific assignment of the hyperparameters and Hamiltonian. We perform these stages classically as quantum computations are costly. However, the final estimation of the quantum algorithm’s cost function is done using a quantum simulator with the optimised hyperparameters.
Machine Learning.
ML, a branch of artificial intelligence (AI), aims to learn from data and generalise models via statistical algorithms (e.g. random forest or linear regression) to identify patterns and make predictions and decisions on new, unseen data. The process typically involves several main phases: data collection, data pre-processing (including data augmentation and mining), model training, and model prediction (or deployment).
In this work: Data consists of Hamiltonians, parameters and the lowest energy levels. We collect data from smaller quantum systems (e.g. 2-qubit, 4-qubit) to build a model for larger systems (20-qubit and above). Data is gathered from online sources or computed via simulations or classical algorithms. The data is then processed and augmented to ensure compatibility with machine learning models, aiming to keep relevant features of the physical system for training. We discuss our choices with respect to data collection, augmentation, and model prediction in Section 5.
Gradient Boosting.
The Gradient Boosting technique [23] is a supervised learning method suitable for regression and classification. It efficiently handles large amounts of data, real numbers, and datasets with many features, as is the case here when representing the Hamiltonian in our dataset, which even with aggressive approximation may result in large data instances. The resultant prediction model allows prediction with large datasets with high precision, avoiding flat predictions caused by scaling issues or oversimplified predictions due to insufficient data relative to the number of features.
XGBoost Library.
We employ XGBoost (eXtreme Gradient Boosting), an open-source Python library of the Gradient Boosting framework [57], as a regressor to predict a response variable, in our case, the lowest energy level of the quantum system (the Hamiltonian). The input features for our model include the Hamiltonian, the number of qubits, and a set of hyperparameters.
4 Quantum Eigensolver Algorithms
During our experiments and development of the solution to the challenge, we used and wrote two implementations of Quantum Eigensolver (QE), ADAPT-QSCI and QCELS, described in Subsection 4.1 and Subsection 4.2. QEs aim to find the nearest approximation of the eigenvalues and eigenvectors of a given system (i.e. the input Hamiltonian). Furthermore, in QC, we aim to determine the ground state energy of the system (the lowest energy level), similarly to the challenge [46].
In our evaluation, we used the implementation of ADAPT-QSCI provided by the challenge scripts [46] and implemented our own QCELS solver. Each implementation can run on a classical simulator (i.e. classic mode) or a NISQ (i.e. QC mode). In classical mode, we are limited by the size of the system and probably can run up to 20-qubit systems at most. In our experiments, we limited our classical mode runs to 16-qubit systems.
Next, we describe the two QE algorithms used in this work.
4.1 ADAPT-QSCI Algorithm
QSCI.
Quantum-selected configuration interaction (QSCI) method is a computational algorithm in quantum chemistry that is used for calculating electronic structure of molecules in an intelligently chosen subspace that makes larger systems feasible to study on modern-day NISQ devices [31]. A full configuration interaction (FCI) requires high computational cost and memory usage that is out of reach for large systems but by using QSCI the computational space can be reduced through selecting only the most important configurations (ways of distributing electrons among molecular orbitals) by a pre-selection algorithm. One such algorithm is ADAPT-QSCI [40].
ADAPT-QSCI.
Adaptive Construction of Input State for Quantum-Selected Configuration Interaction (ADAPT-QSCI) is an iterative algorithm that uses a predetermined pool of single Pauli operators $\text{I\kern-1.49994ptP}=\{P_{1},\dots,P_{T}\}$ that are generators of rotation gates for the input quantum state of QSCI, and selects the best operators from the pool to lower the energy output by QSCI. The operator pools are similar to ADAPT-VQE approaches and could be based on fermionic or qubit excitations. In this method a simple sampling measurement is performed on an input state prepared by a quantum computer - this is the only step that requires quantum computation. The measurement result is used for identifying the most important electron configurations for performing the selected configuration interaction (CI) calculation on classical computers, that is, Hamiltonian diagonalization in the selected $R_{k}$ dimensional subspace $S_{k}=\operatorname{span}\{\ket{r_{1}^{(k)}},\ldots,\ket{r_{R_{k}}^{(k)}}\}$ [40]. QSCI relies on a quantum computer only for generating the electron configurations via sampling, and the subsequent calculations to output the ground-state energy and the pool operator gradients $h_{j}=\braket{c_{k}}{i[H,P_{j}]}{c_{k}}$ are executed on classical computers333|$\ket{c_{k}}$ is a state corresponding to classical vector $c_{k}=\sum_{l=1}^{R_{k}}(c_{k})_{l}\ket{r_{l}^{(k)}}$, and $i[H,P_{j}]$ is calculated through projecting onto the subspace $S_{k}$ and evaluating the expectation classicaly using the classical vector $c_{k}$.[40].. These calculations are made possible on classical machines due to the reduction of the dimensionality of the subspace based on QSCI. The quantum advantage of the ADAPT-QSCI stems from the potential speed up and improved precision of the generating of the electron configurations via sampling.
4.2 Quantum Complex Exponential Least Squares (QCELS) Algorithm
To go beyond NISQ algorithms, where typically the QC is used to simply prepare an ansatz state and then measured in a Pauli basis, we used the Quantum Complex Exponential Least Squares algorithm (QCELS) [17] which uses a controlled time evolution unitary and so introduces a large number of two-qubit gates. However, it does not suffer from the barren plateau problem of VQE and is low enough depth that it will likely be run on early QCs with some error correction, offering a good solution when the number of qubits is too large for NISQ algorithms.
[…]
Figure 1: QCELS circuit [17]
The QCELS algorithm [17] takes a reference state |$\ket{\psi_{0}}$ and evolves it by the time evolution operator $U(t)=e^{-iHt}$ enclosed within a Hadamard test (as depicted by Figure 1). If the reference state is the ground state (|$\ket{\phi_{0}}$) then the resultant expectation values will have a single frequency $Z_{n}=\braket{\phi_{0}}{U(t_{n})}{\phi_{0}}=e^{-iE_{0}t_{n}}$, where $E_{0}$ is the ground state energy. If the reference state, |$\ket{\psi_{0}}$, is not exactly the ground state then the resultant function is $\ Z_{n}\approx\braket{\psi_{0}}{U(t_{n})}{\psi_{0}}=\sum_{i}p_{i}e^{-iE_{i}t_{n}}$ where $p_{i}=|\braket{\phi_{i}}{\psi_{0}}|^{2}$ is the probability of measuring the eigenstate |$\ket{\phi_{i}}$ of the Hamiltonian. Thus, if a reference state with good overlap with the true ground state is known, we can apply the time evolution operator within a Hadamard test $N$ times and fit to the resulting complex exponential. For this challenge, we found that the provided Hamiltonians had an initial Hartree-Fock ground state that retains an overlap with the true ground state - up to $28$ qubits - to extract a good estimate of the energy. The time evolution operator was implemented with a first-order Trotter-Suzuki expansion with Hamiltonian truncated to only the Pauli operators with the largest 200 coefficients to reduce gate count.
Our procedure for fitting a sum of exponentials to the collected data was to initially fit with a single frequency fitting to
$f_{\text{fit}}^{(1)}=r_{1}^{(1)}e^{-i\theta_{1}^{(1)}t}+1-r_{1}^{(1)},$ (1)
before using this frequency as an initial guess and then adding sequentially second and third frequencies,
$f_{\text{fit}}^{(2)}$ $=r_{1}^{(2)}e^{-i\theta_{1}^{(2)}t}+r_{2}^{(2)}e^{-i\theta_{2}^{(2)}t}+1-r_{1}^{(2)}-r_{2}^{(2)}$ (2)
$f_{\text{fit}}^{(3)}$ $=r_{1}^{(3)}e^{-i\theta_{1}^{(3)}t}+r_{2}^{(3)}e^{-i\theta_{2}^{(3)}t}+(1-r_{1}^{(3)}-r_{2}^{(3)})e^{-i\theta_{3}^{(3)}t}$ (3)
which were constrained to be smaller in magnitude than the initial fitted frequency $\theta_{1}^{(1)}$ to ensure that they acted as corrections to the initial fit, for the purpose of the challenge we also added a factor $0.8$ to keep the fitting procedure as safe as possible due to the need for it to be automated. The size of the amplitude $r_{2}^{(2,3)}$ is also constrained to be smaller than $r_{1}^{(1)}$.
The outcome of this procedure when applied to test data collected from an example of the orbital rotated Hubbard Hamiltonian on 28 qubits is given in figure Figure 2, with a sequential improvement towards the correct answer as we increase the number of frequencies in the fit.
(a) $f_{\text{fit}}^{(1)}$
(b) $f_{\text{fit}}^{(2)}$
(c) $f_{\text{fit}}^{(3)}$
Figure 2: QCELS results from sample data of the orbital rotated Hubbard Hamiltonian on 28 qubits (various parameters)
5 Accelerating Quantum Algorithms with Machine Learning
In the previous section, we explored two implementations of the QE algorithms: ADAPT-QSCI and QCELS. Building on these foundations, we now introduce our proposed solution, the prototype of which was submitted to the QAGC challenge [46]. Our approach enhances ADAPT-QSCI and QCELS by integrating machine learning techniques to optimise their hyperparameters. Before presenting our solution, we explain how we gather sufficient data to train an ML model (Subsection 5.1). We then present our solution in detail (Subsection 5.2), followed by implementation details and the constraints set by the QAGC challenge rules (Subsection 5.3). In the evaluation, we demonstrate that quantum algorithms can benefit from incorporating a machine learning model to improve decision-making and overall performance (Section 6).
5.1 Data Augmentation
ML algorithms operate in two primary phases: training and prediction. During the training phase, a sufficiently large dataset is essential for making probabilistic generalisations. In a supervised learning context, this dataset comprises instances where the true output values are known. Once the model is trained, it can make predictions on new, unseen data. The quality and accuracy of the training dataset directly influence the effectiveness of these predictions. Consequently, our initial challenge was to obtain a robust initial dataset.
Acquiring a sufficiently large dataset is a significant challenge. One common solution to this problem is data augmentation, a technique that involves creating additional training data from the existing dataset through various transformations. These transformations can include rotations, scaling, cropping, and other modifications that simulate new data samples, thereby enhancing the diversity and volume of the training dataset. The effectiveness of data augmentation has been extensively studied and validated in numerous literature reviews (e.g., [36, 38, 52, 50, 56]).
The accuracy and generalisation of an ML model are closely linked to the size of the dataset. In the context of quantum simulation data, each data instance encompasses numerous features due to the substantial memory required to represent the Hamiltonian. To generate a sufficiently large dataset for training, we utilise a classical state vector simulator rather than a quantum device, as quantum computing time is prohibitively expensive. This simulator allows us to compute the exact energy levels of smaller Hamiltonians (i.e. fewer than 28 qubits) with high precision. Each data instance in our training dataset encompasses the complete set of information required for the quantum problem, including the Hamiltonian [^4] In practice, we do not include in our data the Hamiltonian as-is but a compressed representation of the Hamiltonian to save memory and avoid overfitting. Compressing a Hamiltonian in the context of QC reduces the complexity of the Hamiltonian operator but keeps the essential physical properties of the system. We compressed by truncating small terms., its exact energy level, and the hyperparameters used to compute this level.
We use this dataset to train our ML model and predict hyperparameters for new Hamiltonians. We anticipate that the generalisation capabilities of deep learning will enable our model to handle larger Hamiltonians and predict hyperparameters for unseen 28-qubit problems effectively. This approach adheres to the challenge rules outlined in Subsection 5.3, as we do not apply classical methods to Hamiltonians of 28 qubits or larger. Further details on the dataset structure and mining processes are discussed in the Data Preparation stage described in Subsection 5.2.
Note on the energy levels predictions: As energy levels are real numbers (i.e., are in a continuous domain), classification might result in precision issues during the prediction phase. Moreover, the exact energy level value is irrelevant as we do not use these values directly. We are interested in knowing whether certain data (Hamiltonian and some quantum-system-related parameters) lead to further minimised answers (the lowest energy level value). This is the relationship between these values. Therefore, we have used a regressor to predict roughly the lowest energy levels.
Figure 3: Classical pre-processing phase - training a regressor on smaller systems
5.2 AccelerQ
In this section, we present AccelerQ, our approach for enhancing quantum algorithm performance through hyperparameter optimisation. We employ a search-based methodology to identify the optimal combination of hyperparameters by leveraging a regressor trained on smaller systems. This regressor estimates system energy, which serves as the fitness function in our optimisation process. Below, we provide an overview of the methods employed, including search-based techniques, the regressor, and its Python implementation.
Figure 3 illustrates our hyperparameter optimisation algorithm for quantum problems, which utilises a regularizing gradient boosting regressor (XGBoost). The process is organised into three key stages: data preparation, model training, and hyperparameter optimisation. This algorithm generates tailored hyperparameter suggestions for executing the QE algorithm, customised for each specific Hamiltonian.
To ensure the generalisability of AccelerQ and its applicability to various QE algorithms (e.g., ADAPT-QSCI or QCELS), we have developed a versatile wrapper. This wrapper encompasses the Hamiltonian, the number of qubits, a parameter indicating whether the computation should be performed classically or on a quantum computer, and the hyperparameters (of both: the QE and ML problems, while we only optimise here those of the QE problem). In our evaluation (see Section 6), we applied AccelerQ to the ADAPT-QSCI and the QCELS algorithms.
Data Preparation.
We add open-source commonly used molecular Hamiltonians (H2O, LiH, BeH2, Hemocyanin [2], and H chain) and combine them with the ones provided by the challenge Hamiltonians of $16$ qubits or less, details in [7]. The wrapper is run in classical mode while giving a randomly generated set of hyperparameters each time, recording the energy level as the score. This process is repeated to produce a data array for training a Regularising Gradient Boosting Regression (XGBoost). The generated data—consisting of a compressed Hamiltonian and hyperparameter vectors (Xs) and their corresponding energy levels (Ys)—is stored for further processing. This stage corresponds to the first of the three steps depicted in Figure 3. We utilise a classical eigensolver during data preparation (Subsection 5.1).
Model Training.
Our algorithm uses an XGBoost model to predict the best hyperparameters. Before training, it ensures that all data vectors are of consistent length by padding them as needed. The padded data array is split into training and testing sets. The model is then trained, tested, and evaluated for performance. The training of the XGBoost model corresponds to the fourth step of Figure 3.
Hyperparameter Optimisation.
With the XGBoost model trained, the algorithm proceeds to predict the optimal hyperparameters for a given Hamiltonian of $28$ qubits (Figure 3, fifth step). It generates a series of hyperparameter vectors and uses the XGBoost model to predict their performance. In each iteration (generation), the best-performing vectors are selected (i.e. predicted to have minimum score value). A crossover operator combines pairs of these vectors, generating new hyperparameters through averaging, extremes, and random values. In the last generation, the vector with the minimum score (predicted energy level) is returned as the optimal set of hyperparameters. Finally, the algorithm runs ADAPT-QSCI with the optimised hyperparameters (Figure 3, sixth and last step).
5.3 Implementation Details
Our implementation is divided into three parts (data collection, training and optimisation) and is written in Python 3.10.12. The rest of the requirements are derived from the rules of the 2024 Quantum Algorithm Grand Challenge (QAGC2024), with one exception: we utilised GPU during the training phase of our models while keeping them relatively small to be able to deploy them on desktop CPU. We used the Python library QURI Parts [^5] https://pypi.org/project/quri-parts/ to simulate and access QC.
Quantum Algorithm Grand Challenge Description.
Participants in the 2024 Quantum Algorithm Grand Challenge (QAGC2024) were tasked with challenging the ADAPT-QSCI Algorithm [40], a quantum-classical hybrid approach designed to compute the ground states and energies of many-body quantum Hamiltonians (Subsection 3.1). The challenge specifically focuses on optimising a quantum algorithm for solving a $28$-qubit system under several constraints (see full list [47]). These constraints included a specific machine specification (CPU, 16 GB RAM), a practical time limit of $6\times 10^{5}$ seconds to ensure all solutions could be evaluated within a reasonable timeframe and a shots limit of $10^{7}$ due to the high cost of quantum computer execution. Participants were required to make essential use of quantum computers, as the goal was a quantum competition, and current classical algorithms do not scale efficiently for systems with $28$ or more qubits.
Furthermore, participants were instructed to output only the energy level of the ground state, calculated using the quantum algorithm of choice. Exact calculations of the ground-state energy using methods like the Bethe ansatz were prohibited, as such methods do not scale well for larger systems ($28$ qubits or more). Additionally, participants were not allowed to hard-code initial parameters or hyperparameters to ensure adaptability to new Hamiltonians. That constraint ensured that the prosposed solution is generic and not tailored for a specific Hamiltonian or subset of problems.
To facilitate the automatic evaluation of all participants’ solutions, the contest organisers provided a specific structure for their responses, limiting participants to modifying a single answer file. While this constraint streamlined the evaluation process, it also introduced certain limitations. Notably, it prevented us from leveraging the full potential of machine learning techniques, as models could not be stored globally and had to be retrained for each evaluation. Although this approach was effective for the contest evaluation, it is not ideal for software engineering practices that prioritise modularity. A more effective strategy would be to divide the problem into distinct stages, each with its own constraints. Constraints provided by the contest organisers are particularly relevant for the prediction phase but may not be suitable for the training phase. Recognising these limitations, in future iterations of our solution we plan to adopt further a more flexible approach, allowing for greater use of advanced methods and better adaptability, while respecting the original framework and intentions set by the challenge organisers.
6 Evaluation
We evaluated AccelerQ for its ability to further optimise ADAPT-QSCI and QCELS algorithms by suggesting better hyperparameters tailored per system.
6.1 Methodolgy
We demonstrated our idea on the optimisation of hyperparameters of two quantum eigensolver algorithms. We used 4- to 16-qubit systems to train the models. We deployed the models on 20-, 24- and 28-qubit systems. We aim to understand:
RQ1: To what extent can AccelerQ’s optimisation of hyperparameters alone accelerate and improve the efficiency and accuracy of Quantum Eigensolver (QE) algorithms on NISQ devices in terms of runtime, error, and system size?
Given that we extract a model from smaller systems up to 16 qubits, we wish to assess the scalability of these models and to test if we have reduced performance as the system gets larger, that is:
RQ2: How scalable are machine learning-predicted hyperparameters learned on smaller systems when applied to QE algorithms for Hamiltonian systems with increasing qubit number and complexity?
In our evaluation, we applied our methodology to 16 larger systems of 20-, 24- and 28-qubit with known lowest energy levels, using the two QE implementations: ADAPT-QSCI and QCELS (Section 4) to answer RQ1 and RQ2.
We constructed two models—one for each implementation—using data extracted classically (Section 4) from up to 16-qubit systems. The data extraction and model training were general. The resultant model aimed to predict the optimal hyperparameters for its QE implementation and a system (Hamiltonian). The optimisation process took an implementation, its trained model and a system (a problem to solve) as input and returned the optimal hyperparameters for this setup.
Baseline and Parametrisation.
Our evaluation baseline is the results obtained with an implementation’s default hyperparameters, fixed across all systems in the evaluation. The hyperparameters were specific per implementation, were optimised, and their performance was compared against the baseline. Apart from the hyperparameters, ADAPT-QSCI and QCELS implementations were provided with ham and number_qubits, the input system and the required number of qubits for the corresponding Hamiltonian, and the flag is_classical being set to True during the data collection phase, and either True or False otherwise (classical or QC). We did not optimise this part.
ADAPT-QSCI. The ADAPT-QSCI implementation [40] includes several hyperparameters that control its operation. These hyperparameters have default values, which we used as our baseline for comparison. num_pickup (default: 100) and coeff_cutoff (default: 0.001) parameters control the terms retained or removed from the Hamiltonian compressed representation. self_selection indicates whether self-selection is enabled (default: False), which forces the algorithm to work in subspace. iter_max is the maximum total number of iterations (default: 100). sampling_shots is the number of sampling shots for measurements per iteration (default: $10^{5}$). atol is the absolute tolerance for convergence criteria (default: $61e-$ 6). final_sampling_shots_coeff is the number of shots for the calculations if the same operator appears twice or the operator parameter is close to zero (default: 5). num_precise_gradient is the number of operators from the pool to calculate gradient more precisely (default: 128). max_num_converged specifies the maximum number of iterations within atol needed for a solution to be considered converged (default: 2). reset_ignored_inx_mode specifies after how many iterations previously used operators can be reused (default: 0). These default values are the baseline for configuring the ADAPT-QSCI implementation, but these can be overridden (e.g.) when using AccelerQ suggestions tailored per system.
QCELS. The QCELS hyperparameters are a different set of arguments. This includes the following: ham_terms is the number of individual terms that are retained in the Hamiltonian after truncation. The Hamiltonian is transformed to a qubit Hamiltonian formed from a linear combination of Pauli strings, the truncation then retains the largest ham_terms of these and the rest are discarded. The default value is set to be 200 and we search for its optimal value between random.randint(50, 1000). ham_cutoff is the same as coeff_cutoff in ADAPT-QSCI implementation, setting a minimum value for the retained coefficients, discarding all terms in the Hamiltonian with coefficients lower than this threshold. delta_t is the time step for the simulation or evolution of the system (the default is 0.03; delta_t = random.uniform(1e-3, 0.3)), the expectation value of the time evolution operator is calculated at a set of times starting from zero and separated by this value. n_Z is the total number of points used in fitting the time evolution and so n_Z11-1- 1 is the total number of points at which the expectation value is evaluated. The default value is 10; n_Z = random.randint(5, 25). alpha is a scalar parameter that determines how much smaller the parameters $r_{2}^{(2)}$, $\theta_{2}^{(2)}$ and $r_{2}^{(3)}$, $\theta_{2}^{(3)}$, $\theta_{3}^{(3)}$ are forced to be than (default is $0.8$ ;alpha = random.uniform(0.5, 1)).
Using the above, we defined different types of hyperparameter vectors per implementation. The input vectors (’X’s) were normalised to the size of vectors in 28-qubit systems and included the implementation hyperparameters and the system’s Hamiltonian. The Hamiltonians were compressed by removing small elements (abs(0.05) and below). The target values (Y’s) represent the predicted lowest energy levels. Note: since we utilised the classical mode when extracting data from smaller systems, these are rough approximations, not the true values.
6.2 Experimental Setup
We extracted 66 files for the training phase using classical wrappers (i.e. 33 for ADAPT-QSCI and 33 for QCELS). We used open-source molecular Hamiltonians–smaller systems prepared as discussed in the data preparation paragraph in Subsection 5.2– and the challenge Hamiltonians–the Hamiltonians of 04 and 12 qubits for seeds __00 to __04 from [44].
We trained the ADAPT-QSCI with 4760 records (small dataset) and QCELS with 14510 records (medium dataset), out of which 400 and 5550 records (respectively) with the challenge Hamiltonians. Per system sizes of 4-, 6-, 7-, 8-, 10-, 12-, 14-, and 16-qubit, we utilised 60, 750, 500, 500, 1500, 450, 800 and 200 records for ADAPT-QSCI hyperparameters, and 60, 750, 500, 500, 1000, 6600, 2100 and 3000 for QCELS hyperparameters, respectively. The QCELS implementation ran faster than ADAPT-QSCI, allowing us to extract a medium-sized dataset for QCELS. The models were trained using XGBRegressor [^6] https://xgboost.readthedocs.io/en/latest/python/python_api.html#xgboost.XGBRegressor, version 2.1.1 of the XGBoost library, on a single GPU core, NVIDIA 12GB PCI P100 GPU, 12 GB VRAM, running Ubuntu 22.04.4 LTS. We deployed the model on an ARM M2 Mac running Ubuntu 22.04.4 with 8 GB RAM.
We evaluated on a quantum simulator the predictions for several Hamiltonians of 20-, 24-, and 28-qubit systems with a known real answer (the challenge Hamiltonians were 20 and 28 qubits for seeds __00 to __04 [44]; the rest of the seeds were open-source molecular Hamiltonians Subsection 5.2). We ran the simulations on a single virtual machine with 16 virtual CPU cores and 32 GB RAM running Ubuntu 20.04.2 LTS x86_64. The host had a single AMD EPYC 7313P CPU (single socket, 3.0 GHz, 16 cores, 2 threads per CPU).
6.3 Results
The models for ADAPT-QSCI and QCELS were trained with dataset sizes of 757 MB and 4868 MB, respectively. The model sizes were 1.1 MB for ADAPT-QSCI and 2.86 MB for QCELS.
We first evaluated the model performance by calculating the mean absolute error (MAE) and the mean squared error (MSE) of two testing sets. We sampled 10% of data from each system (first testing set) and split the remaining data into training and (second) testing sets at an 80% - 20% ratio. We had no validation stage due to (1) the limited number of systems with known Y’s value (typically sourced from physics publications) and (2) only classically estimating the Ys for a hyperparameters assignment. Our primary focus is evaluating the algorithm AccelerQ, not the ML component.
For ADAPT-QSCI, the MAE and MSE were 6.28776 and 129.512 (for the 10% set) and 6.37912 and 122.659 (for the 20% set), and for QCELS, these were 9.05254, 258.031, 8.25527 and 145.092. These results indicate the need for a custom loss function, some feature weighting, and a validation stage with larger datasets, which we leave for future work.
Table 1: Predicted optimal hyperparameters (A-J col.) with the first row stating the default values for ADAPT-QSCI implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: num_pickup, B: coeff_cutoff, C: self_selection, D: iter_max, E: sampling_shots, F: atol, G: final_sampling_shots_coeff, H: num_precise_gradient, I: max_num_converged, and J: reset_ignored_inx_mode)
System Size A B C D E F G H I J
Default - 100 0.001 0 100 100000 1.00E-06 5 128 2 0
20qubits_00 63636 985 7.64647e-03 0 344658 49458 5.67168e-05 3 77 2 1
20qubits_01 54717 807 9.41494e-03 1 382109 964853 9.00118e-05 6 79 2 64
20qubits_02 54723 934 4.56117e-03 0 478268 875460 4.27314e-05 8 294 3 83
20qubits_03 54710 551 1.13139e-03 1 802530 106374 7.23056e-05 1 161 2 79
20qubits_04 63636 182 4.06265e-03 0 398085 61950 7.69414e-05 7 194 4 11
20qubits_05 46 593 8.46933e-04 1 278431 148423 6.66628e-05 7 158 4 58
24qubits_05 56 67 7.77741e-03 1 655715 676443 3.37411e-05 8 176 2 89
24qubits_06 56 292 4.00805e-03 0 762863 603105 6.33079e-05 3 159 4 82
24qubits_07 56 80 4.06535e-03 0 659716 711486 9.45728e-05 4 37 3 0
24qubits_08 56 215 6.59156e-03 1 885413 291401 1.74522e-05 1 247 4 81
24qubits_09 56 716 7.29064e-03 0 401350 929361 6.82934e-05 5 257 2 65
28qubits_01 98700 1000 6.58567e-03 0 615787 300938 1.04770e-05 8 224 4 98
28qubits_00 98700 197 1.42217e-03 1 460054 209518 7.62522e-05 7 287 3 54
28qubits_02 98700 531 5.88032e-03 1 402352 258683 4.81220e-06 8 277 2 47
28qubits_03 98700 629 4.35568e-03 0 429007 82966 3.89304e-05 8 239 4 8
28qubits_04 98700 698 7.96341e-03 0 221061 633412 7.86532e-05 4 68 4 89
Table 2: Predicted optimal hyperparameters (A-E cols.) with the first row stating the default values for QCELS implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: delta_t, B: n_Z, C: ham_terms, D: ham_cutoff, and E: alpha)
System Size A B C D E
Default - 0.03 10 200 1e-9 0.8
20qubits_00 63636 1.37929e-01 14 877 6.77275e-03 9.76144e-01
20qubits_01 54717 2.87833e-01 8 69 9.23310e-04 5.26429e-01
20qubits_02 54723 1.05366e-01 10 989 4.31098e-04 8.77018e-01
20qubits_03 54710 1.04235e-01 16 68 6.91101e-03 5.12876e-01
20qubits_04 63636 1.31645e-01 19 971 2.10235e-03 7.76523e-01
20qubits_05 46 1.98379e-01 7 187 8.34526e-03 5.33536e-01
24qubits_05 56 2.08699e-01 24 933 5.75035e-03 6.20185e-01
24qubits_06 56 1.75486e-01 10 398 7.28888e-03 9.72887e-01
24qubits_07 56 1.80166e-01 11 286 1.06348e-03 9.75924e-01
24qubits_08 56 2.72722e-01 22 453 6.67156e-03 8.29813e-01
24qubits_09 56 2.09069e-01 24 650 3.99377e-03 5.20032e-01
28qubits_00 98700 7.00174e-03 7 310 6.71154e-03 6.32763e-01
28qubits_01 98700 1.99713e-01 23 309 8.39957e-03 5.54053e-01
28qubits_02 98700 5.46760e-02 12 997 1.22706e-03 5.60243e-01
28qubits_03 98700 4.23353e-02 6 553 8.40543e-03 8.66162e-01
28qubits_04 98700 2.33005e-01 25 280 1.30071e-04 5.98526e-01
Table 3: Hamiltonian sizes (Size col.) and the size after compression used when querying the ML Models (Comp. col.). The column Comp. represents the size of the system as utilised in the ML model. We marked with an asterisk the open-source molecular Hamiltonians.
System Size Comp.
20qubits_00 63636 91
20qubits_01 54717 287
20qubits_02 54723 445
20qubits_03 54710 338
20qubits_04 63636 65
20qubits_05(*) 46 46
System Size Comp.
24qubits_05(*) 56 48
24qubits_06(*) 56 46
24qubits_07(*) 56 50
24qubits_08(*) 56 51
24qubits_09(*) 56 47
System Size Comp.
28qubits_00 98700 167
28qubits_01 98700 158
28qubits_02 98700 98
28qubits_03 98700 144
28qubits_04 98700 213
Optimal Hyperparameters Prediction Results.
Table 1 and Table 2 show the optimal hyperparameters found by AccelerQ with each Hamiltonian (System col.) employing the two models above. Size col. is the number of terms in the system. The first row shows the default values of the implementation. Table 1 shows the ADAPT-QSCI’s optimal hyperparameters, with col. A-J represents the predicted optimal hyperparameters values. Similarly, Table 2 presents the optimal hyperparameters for the QCELS implementation in col. A-E. The value presented in Table 1 for the predicted iter_max was capped during execution [^7] The platform automatically stopped the computation once the maximum number of shots, 10 000 0001000000010\,000\,00010 000 000, was reached. to ensure that we do not have unlimited resources. In practice, the iter_max has the same limit as with the default parameters (i.e. iter_max=1e7/sampling_shots; col. E and col. D). Table 3 shows the size (Size col.) and the compressed size (Comp. col.) of the Hamiltonians (System col.) used when querying the ML model during prediction, marking with an asterisk the open-source molecular Hamiltonians. Different cutoffs and compression rates were used for ML part than in the QE implementations. Consequentially, the sizes recorded in Table 3 are true only for the ML model.
In Table 3, the number of terms in the open-source molecular Hamiltonians systems ($\leq 46$) is significantly smaller than the challenge Hamiltonians ones ($\geq 54k$), even after compression, the open-source molecular Hamiltonians systems was still much smaller on average. This suggests that these systems may differ to some extent.
The results in Table 1 suggest that the optimisation opted to (sometimes) lower the performance of each iteration, resulting in less precise results per iteration but using more iterations overall by reducing (sometimes) sampling shots and increasing the atol, coeff_cutoff and iter_max. However, when the optimisation predicted both iter_max and sampling_shots at maximum values, ignoring their correlation, the execution was capped, resulting in fewer iterations in practice.
The results in Table 2 show that the optimisation favoured increasing the number of large coefficients (except for 20qubits_01 and 20qubits_03) using larger ham_cutoff values and hence leaving fewer small coefficients. Further, the optimisation selected higher values for n_z and delta_t, with no clear preference for alpha.
Table 4: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
System Value Algorithm Error (%) Task Score Aprox. Runtime Itr.
20qubits_00 -22.046059902 ADAPT-QSCI, Opt. 2.46 -21.503637869562557 15777 s 159
ADAPT-QSCI, Default 4.52 -21.05032811997207 8211 s 97
QCELS, Opt. 2.57 -21.48036468528949 14991 s 157
QCELS, Default 4.47 -21.0598063085247 7936 s 92
20qubits_01 -22.046059902 ADAPT-QSCI, Opt. 6.26 -20.66624780754983 4476 s 11
ADAPT-QSCI, Default 4.49 -21.055920933889066 7463 s 84
QCELS, Opt. 6.30 -20.655947060040507 4327 s 11
QCELS, Default 4.30 -21.097506543499456 7486 s 82
20qubits_02 -22.046059902 ADAPT-QSCI, Opt. 5.85 -20.75547153586126 4941 s 12
ADAPT-QSCI, Default 4.55 -21.042520869027147 6370 s 65
QCELS, Opt. 5.85 -20.756855659992507 4789 s 12
QCELS, Default 4.62 -21.02776101198928 6152 s 65
20qubits_03 -22.046059902 ADAPT-QSCI, Opt. 3.37 -21.302881787984727 8631 s 95
ADAPT-QSCI, Default 4.63 -21.024636461355875 6682 s 77
QCELS, Opt. 3.27 -21.32597699721156 8575 s 95
QCELS, Default 4.59 -21.03455240312227 6293 s 66
20qubits_04 -22.046059902 ADAPT-QSCI, Opt. 3.31 -21.316521673460805 11004 s 121
ADAPT-QSCI, Default 4.64 -21.024150729181112 7595 s 88
QCELS, Opt. 3.34 -21.308808719093236 9807 s 108
QCELS, Default 4.57 -21.03798918026242 7039 s 83
28qubits_00 -30.748822808 ADAPT-QSCI, Opt. 6.35 -28.7974854444735 10812 s 48
ADAPT-QSCI, Default 6.42 -28.77456507714568 10345 s 61
QCELS, Opt. 6.35 -28.794794630409676 11076 s 48
QCELS, Default 6.40 -28.779806038153325 10989 s 65
28qubits_01 -30.748822808 ADAPT-QSCI, Opt. 6.39 -28.783766574436143 10024 s 34
ADAPT-QSCI, Default 6.47 -28.758306955348417 11789 s 77
QCELS, Opt. 6.41 -28.777776815200454 10096 s 34
QCELS, Default 6.46 -28.763094962187317 11249 s 74
28qubits_02 -30.748822808 ADAPT-QSCI, Opt. 6.41 -28.778922785865355 11373 s 39
ADAPT-QSCI, Default 6.68 -28.69374047078425 9479 s 53
QCELS, Opt. 6.39 -28.7835710030780 10666 s 39
QCELS, Default 6.66 -28.70067301821086 9509 s 57
28qubits_03 -30.748822808 ADAPT-QSCI, Opt. 6.25 -28.827497968717516 11248 s 51
ADAPT-QSCI, Default 6.47 -28.75988277903682 12548 s 85
QCELS, Opt. 6.18 -28.848400001238723 11644 s 53
QCELS, Default 6.66 -28.699830076240456 11168 s 70
28qubits_04 -30.748822808 ADAPT-QSCI, Opt. 6.79 -28.662288396044573 6757 s 16
ADAPT-QSCI, Default 6.77 -28.667008949938953 8603 s 43
QCELS, Opt. 6.78 -28.66351868143365 7463 s 16
QCELS, Default 6.70 -28.689855023621764 5909 s 45
Table 5: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
System Algorithm Task Score Aprox. Runtime Itr.
20qubits05 ADAPT-QSCI, Opt. -13.120566186315502 3568 s 27
ADAPT-QSCI, Default -14.000000000000009 686 s 9
QCELS, Opt. -12.968777862300158 3719 s 26
QCELS, Default -14.000000000000004 681 s 9
24qubits05 ADAPT-QSCI, Opt. -9.034894851637292 3029 s 11
ADAPT-QSCI, Default -10.99319108738972 7053 s 35
QCELS, Opt. -8.622105677825234 4741 s 8
QCELS, Default -9.039194172208862 5367 s 35
24qubits06 ADAPT-QSCI, Opt. -9.460283005059706 4718 s 8
ADAPT-QSCI, Default -9.460283005059702 644 s 6
QCELS, Opt. -9.460283005059704 4467 s 8
QCELS, Default -9.4602830050597 663 s 6
24qubits07 ADAPT-QSCI, Opt. -7.34876884915856 4637 s 12
ADAPT-QSCI, Default -10.400021274590904 8117 s 37
QCELS, Opt. -7.932213481832702 3977 s 10
QCELS, Default -9.973622581884937 5117 s 36
24qubits08 ADAPT-QSCI, Opt. -6.274613702752347 4560 s 35
ADAPT-QSCI, Default -6.034877031453911 1516 s 17
QCELS, Opt. -5.598926935684615 4779 s 35
QCELS, Default -6.034877031453906 1585 s 17
24qubits09 ADAPT-QSCI, Opt. -13.136457290516125 4660 s 7
ADAPT-QSCI, Default -13.136457290516129 649 s 7
QCELS, Opt. -13.13645729051612 3941 s 6
QCELS, Default -13.13645729051613 686 s 7
Results of Execution with Different Hyperparameters.
Table 4 and Table 5 summarise the results of executing ADAPT-QSCI and QCELS on a quantum simulator with the default and optimal hyperparameters in Table 1 and Table 2. We evaluated predictions on several Hamiltonians of 20-, 24-, and 28-qubit systems (System col.) with a known solution (Value col.). Table 4 presents the result obtained with the challenge Hamiltonians, while Table 5 is the results using the open-source molecular Hamiltonians. Both tables present, for each algorithm and set of hyperparameters (Algorithm col.), the task score (Task Score col.), the approximate run time (Approx. Runtime col.), and the number of iterations completed (Itr. col.). For Table 4, the true results (Value col.) are known [46], and we computed the relative error (Error (%) col.) per experiment. We ran each experiment several times and selected the best result.
In Table 4, the optimised hyperparameters tended to achieve better results than the default. The optimised ADAPT-QSCI found the best solution for 4 systems, optimised QCELS for 3, default QCELS for 2, and default ADAPT-QSCI for 1 out of 10 tested systems. For 20-qubit systems, optimised ADAPT-QSCI had the lowest error at 4.25%, followed by optimised QCELS at 4.27%. Default QCELS and ADAPT-QSCI had average error rates of 4.51% and 4.57%, respectively. For 28-qubit systems, optimised QCELS had the lowest error at 6.42%, followed by optimised ADAPT-QSCI at 6.44%, default ADAPT-QSCI at 6.56%, and default QCELS at 6.58%. Table 5 shows that default parameters performed better. The default ADAPT-QSCI found the best solution for 3 systems, followed by optimised ADAPT-QSCI for 2, and default QCELS for 1 out of 6 tested systems.
Answer to RQ1.
The results indicate a limited ability to improve solely through hyperparameter optimisation, though the optimised hyperparameters performed better in Table 4. A more refined model based on Hamiltonian characteristics is required to properly assess the optimisation of hyperparameters’ impact on accelerating and improving QE efficiency and accuracy.
In detail: (20-qubit) Both implementations, with default or optimised hyperparameters, had similar iteration counts with the default setup consuming less time on average. However, optimised setups performed worse when the number of iterations was too low. (24-qubit) Time consumption increased for both ADAPT-QSCI and QCELS with optimised hyperparameters and used fewer iterations when the error percentage was higher than in the default setup. (28-qubit) The number of iterations was halved on average, with optimised setups requiring 38 compared to 62-64 iterations with default setups. For all systems, optimised hyperparameters used 42-44 iterations versus 53-56 for the default setups, with some increase in time consumption for the optimised setup. 9 systems performed better with optimised setups, while 7 performed better with the default setups, mostly from Table 5.
Answer to RQ2.
Our understanding of AccelerQ’s scalability remains somewhat limited: 80% of the 28-qubit systems’ scores improved with optimised setup, while only 45% of the 20- and 24-qubit systems (5 out of 11) showed improvement. This suggests other factors, like Hamiltonian nature, hyperparameter types, padding, or even data proportions (per source of Hamiltonians), affect scalability and require further investigation.
In detail: Table 4 shows that optimised hyperparameters performed better for the challenge Hamiltonians. We observed the opposite for the open-source molecular Hamiltonians Table 5, particularly for optimised QCELS. The challenge Hamiltonian systems contributed 8.4% of ADAPT-QSCI’s and 37.9% of QCELS’s training data. There is also a correlation between the models: when QCELS with optimised hyperparameters outperformed the default setup, ADAPT-QSCI typically did as well, and vice versa. This suggests that the data each type of Hamiltonian contributed might not be the primary factor. Lastly, the open-source molecular Hamiltonians systems had fewer terms used in the ML phase in comparison to the challenge Hamiltonians, which may affect the model’s ability to generalise [^8] Open-source molecular Hamiltonians averaged 48.2 elements, compared to 200.6 for the challenge Hamiltonians, Table 3..
7 Conclusion
In this paper, we presented interdisciplinary work that merges software engineering and machine learning paradigms to enhance quantum algorithms’ performance on NISQ hardware. This idea explored using Hybrid Quantum-Classical Systems as a promising way to utilise current NISQ hardware. While classical computers are essential to optimise noisy quantum hardware, the workload balance should still keep the quantum device at the core of the computation for such systems to stay relevant. The limit between optimising the quantum circuit outcome and bypassing the hardware limitations with classical means is difficult to define; it is however important to keep this balance in mind when developing hybrid systems.
We designed a new framework, implemented as a prototype tool, AccelerQ, to predict quantum algorithm (near to) optimal hyperparameters. We evaluated AccelerQ on two implementations, training and deploying relatively small-scale models to improve performance by suggesting better hyperparameters. Our results raised the possibility that the model’s ability to predict optimal hyperparameters depends on the Hamiltonian characteristics rather than solely on a specific implementation (ADAPT-QSCI or QCELS).
Code and Data Availability. The code, the training data, the models and the results are available as open-source at [7, 8].
Acknowledgments. We thank CloudLab [18] for the platform and infrastructure support that enabled the experiment resulting in the construction of two models for accelerating quantum algorithms. Sophie Fortz, Connor Lenihan and Avner Bensoussan are partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1. Sophie Fortz and Avner Bensoussan are partially funded by the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009. Sophie is partially funded by Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1 and by the QAssure project from Innovate UK. Also King’s Quantum grants provided by King’s College London are gratefully acknowledged.

### v2

@@
 We evaluated our preliminary approach on 20-, 24- and 28-qubit systems by optimising the Eigensolver’s hyperparameters.
-These models predict hyperparameter values, leading to a 0.13%-0.15% reduction in error when tested on 28-qubit systems.
+These models predict hyperparameter values, leading to a 0.12% reduction in error when tested on 28-qubit systems.
 However, due to inconclusive results with 20- and 24-qubit systems, we suggest further examination of the training data based on Hamiltonian characteristics.
@@
 In this interim time of intensive hardware development, robust and innovative quantum algorithms [39] provide a way to tackle the limitations of the current quantum processors.
-One of the most promising applications of QC is in the field of quantum chemistry for molecular simulations, structure design, drug design and more [37, 33, 35].
+One of the most promising applications of QC is in the field of quantum chemistry for molecular simulations, structure design, drug design, and more [45, 38, 43].
 A prominent example of a class of quantum algorithms that is rapidly evolving in the field of quantum chemistry is the Variational Quantum Eigensolver (VQE) [67], which aims at estimating the ground state energy of Hamiltonians.
@@
 Our approach focuses on applying machine learning algorithms to explore and optimise quantum algorithms to achieve potential advancements in quantum eigensolving and improve the performance and accuracy of quantum algorithms for larger system setups that are currently too complex to tackle.
-Our first attempt at the problem was to find a methodology to select the best hyperparameters for the ADAPT-QSCI quantum algorithm [32] that minimise the runtime and the final result for the ground state energy through using machine learning techniques trained on smaller systems [13].
+Our first attempt at the problem was to find a methodology to select the best hyperparameters for the ADAPT-QSCI quantum algorithm [36] that minimises the runtime and the final result for the ground state energy through using machine learning techniques trained on smaller systems [10].
 This method is not specifically tailored to ADAPT-QSCI and can be applied to any Quantum Eigensolver-based algorithm.
@@
 We investigated the suitability and limitations of the QCELS algorithm when applied to large systems [40].
-In the evaluation, we have assessed their performance and potential benefits for solving 16 quantum systems of 20-, 24- and 28-qubits, showing limited minor improvement for 20- and 28-qubit systems.
+In the evaluation, we have assessed their performance and potential benefits for solving 16 quantum systems of 20-, 24- and 28-qubits, showing improvement for some of the 20- and 28-qubit systems.
 The rest of the paper is structured as follows.
-Section 2 lists related work in the field and Section 3 focuses on the background of the quantum computing algorithms that we have investigated, software engineering optimisation methods and machine learning, followed by a deeper dive into the methodologies of quantum eigensolver algorithms in Section 4, and machine learning applications to accelerate quantum algorithms in Section 5.
+Section 2 lists related work in the fields of quantum computing and machine learning for quantum.
+Section 3 focuses on the background of the quantum computing algorithms that we have investigated, software engineering optimisation methods and machine learning, followed by a deeper dive into the methodologies of quantum eigensolver algorithms in Section 4, and machine learning applications to accelerate quantum algorithms in Section 5.
 In Section 6, we present the experiment’s methodology and our results.
@@
 the number of shots of $10^{7}$, timeout of $6\times 10^{5}$ seconds and limited input data), that can output the closest answer to the actual ground state energy.
-Previous year’s challenge [45] had the same aim but was based on a much smaller 8-qubit system.
-The winning algorithm was the company’s proposition later published as Quantum-Selected Configuration Interaction (QSCI) [32].
+The 2023 challenge [57] had the same aim but was based on a much smaller 8-qubit system.
+The winning algorithm was the company’s proposition, later published as Quantum-Selected Configuration Interaction (QSCI) [36].
 2 Related Work
-Quantum Computing.
-Quantum Computing (QC) was conceptualised initially to simulate quantum mechanics using computers “built of quantum mechanical elements which obey [the] quantum mechanical law” [22].
+Challenges in Quantum Computing.
+Quantum Computing (QC) was conceptualised initially to simulate quantum mechanics using computers “built of quantum mechanical elements which obey quantum mechanical law” [26].
 Later, it was found that QC could have several potential applications and offer significant speed-up over classical computing [6, 5, 17, 7, 31, 13].
 In 1994, Shor’s proposal of a polynomial-time algorithm for prime factorization and discrete logarithms on a quantum computer raised enormous interest due to its potential threat to modern RSA cryptosystems [63].
-Soon after, Grover introduced a fast database search on quantum computers that promised quadratic speed-up over the best classical algorithm [27].
+Soon after, Grover introduced a fast database search on quantum computers [31] that promised quadratic speed-up over the best classical algorithm.
 The resulting potential speed-up is often referred to as “quantum supremacy” [3].
-Several studies have applied software engineering techniques to optimise quantum computing [24, 25], while testing and debugging approaches have been found to be beneficial in quantum software development [39, 59].
-NISQ.
+Several studies apply software engineering techniques to optimise quantum computing [29, 28].
+Noticeably, testing [48, 65], debugging [51, 62, 46], verification [76, 41] approaches, and efficient synthesis techniques [35, 70, 52] have been found to be beneficial in quantum software development [48, 65, 51, 76].
 Demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where quantum devices would solve real-life calculations.
-Although quantum supremacy seems difficult to achieve soon, NISQ algorithms are a prominent example that hybrid systems combining small quantum circuits with classical computations could present some computational advantages, i.e., a quantum advantage [43].
+Although quantum supremacy seems difficult to achieve soon, NISQ algorithms—Imperfect hardware is often called Noisy Intermediate-Scale Quantum (NISQ) devices—are a prominent example that hybrid systems combining small quantum circuits with classical computations could present some computational advantages, i.e., a quantum advantage [55].
 Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era [55].
 Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop [67].
-Machine Learning.
-Machine Learning (ML) algorithms are increasingly used to improve and automate software engineering tasks [55, 58, 28, 48, 19], especially after the advent of Large Language Models (LLM), with common applications in software engineering, including optimisation, code generation, bug detection and automated testing [20, 53, 12, 9, 15, 30].
-Furthermore, connections between Machine Learning and QC have been broadly explored, both to optimise Machine Learning with QC and to optimise QC with Machine Learning [21, 41, 54].
-As we aim to optimise the quantum algorithm hyper-parameters classically, we discuss how we apply machine learning to enhance the software engineering aspects of our approach in Section 5.
+Other algorithms use classical optimisation to optimise quantum calculations, such as QCELS [20, 21] that uses a fitting procedure to optimise quantum calculations.
+Because of their prominent role in modern Quantum Computing research and industrial applications, we chose to focus our study on Variational and QCELS Algorithms.
+Machine Learning for Quantum Software Engineering.
+Machine learning algorithms are increasingly used to improve and automate software engineering tasks [74, 77, 32, 61, 23], especially after the advent of Large Language Models (LLM), with common applications in software engineering, including optimisation, code generation, bug detection and automated testing [24, 72, 14, 11, 16, 34].
+In Section 5, we suggest applying these ideas in the context of quantum algorithms, accelerating their performances We apply a search-based process to identify the optimal combination of the hyperparameters (global minima) using our regressor trained on smaller systems to estimate the system energy as the fitness function.
+We give here an overview of the methods used: search-based methods, the regressor and its Python implementation.
+[25, 73, 53].
+Although existing applications of Machine Learning in Quantum Computing have been investigated [60], it remains a very young and open field of research.
+While Quantum Machine Learning offers potential speedups in Machine Learning tasks [69], very recent work also demonstrates promising results in applying Machine Learning to Optimise Quantum Computations [15, 42].
 3 Background
@@
   3. 3.
-    Convergence check and output The optimisation process continues until a convergence criterion is met, indicating that further iterations are unlikely to significantly improve the solution.
+    Convergence check and output.
+The optimisation process continues until a convergence criterion is met, indicating that further iterations are unlikely to significantly improve the solution.
 This convergence check ensures that the algorithm has reached a stable and potentially optimal solution.
@@
 ADAPT-VQEs are found to have more precise results in some applications, at a cost of a higher computational load.
-3.2 Use of AI with GI and Search-based Optimisation
+3.2 Machine Learning and Software Engineering Methods
 In this work, we integrate ML with search-based methods.
 We first use predictions on smaller systems, utilising the cost function values of a Hamiltonian of smaller systems and randomly generated hyperparameters of the quantum algorithm to train a gradient boosting model to predict the cost function value.
-We then use this model to explore and optimise hyperparameters for a larger Hamiltonian system, which is costly to solve classically, via a crossover operator.
-Search-based Optimisation and Genetic Improvement (GI).
+We then use this model to explore and optimise hyperparameters for a larger Hamiltonian system.
+Search-based Software Engineering.
 Search-based methods [33] have also been used for optimisation tasks to find the best possible subset of requirements.
@@
 Starting with a set of candidate solutions, they are evaluated using a fitness function.
-GI then diversifies solutions over iterations through mutation operations like bit flip (on a single solution) or crossover (on multiple solutions) to potentially generate better-performing offspring.
+In this work, we utilise the genetic algorithms (GA) approach.
+GA diversifies solutions over iterations through mutation operations like bit flip (on a single solution) or crossover (on multiple solutions) to potentially generate better-performing offspring.
 The method can occasionally minimise the population or inject noise to avoid converging to local minima (or maxima, depending on the problem).
@@
 In this work: Data consists of Hamiltonians, parameters and the lowest energy levels.
-We collect data from smaller quantum systems (e.g.
-2-qubit, 4-qubit) to build a model for larger systems (20-qubit and above).
+We collect data from smaller quantum systems (16 qubits and below) to build a model for larger systems (20 qubits and above).
 Data is gathered from online sources or computed via simulations or classical algorithms.
@@
 Gradient Boosting.
-The Gradient Boosting technique [23] is a supervised learning method suitable for regression and classification.
-It efficiently handles large amounts of data, real numbers, and datasets with many features, as is the case here when representing the Hamiltonian in our dataset, which even with aggressive approximation may result in large data instances.
+The Gradient Boosting technique [27] is a supervised learning method suitable for regression and classification. Supervised machine learning utilises labelled data to train predictive models [37].
+Gradient boosting efficiently handles large amounts of data, real numbers, and datasets with many features, as is the case here when representing the Hamiltonian in our dataset, which, even with aggressive approximation, may result in large data instances.
 The resultant prediction model allows prediction with large datasets with high precision, avoiding flat predictions caused by scaling issues or oversimplified predictions due to insufficient data relative to the number of features.
@@
 In our evaluation, we used the implementation of ADAPT-QSCI provided by the challenge scripts [58] and implemented our own QCELS solver.
-Each implementation can run on a classical simulator (i.e.
-classic mode) or a NISQ (i.e.
-QC mode).
+Each implementation can run on a classical statevector (exact) simulator (i.e.
+classic mode) or on a matrix product state (MPS) simulator (i.e.
+QC mode) [79, 4, 78, 50], provided as part of the challenge, which is suitable for simulating large system sizes by enabling a trade-off between speed and accuracy.
 In classical mode, we are limited by the size of the system and probably can run up to 20-qubit systems at most.
@@
 Quantum-selected configuration interaction (QSCI) method is a computational algorithm in quantum chemistry that is used for calculating electronic structure of molecules in an intelligently chosen subspace that makes larger systems feasible to study on modern-day NISQ devices [36].
-A full configuration interaction (FCI) requires high computational cost and memory usage that is out of reach for large systems but by using QSCI the computational space can be reduced through selecting only the most important configurations (ways of distributing electrons among molecular orbitals) by a pre-selection algorithm.
+A full configuration interaction requires high computational cost and memory usage that is out of reach for large systems but by using QSCI the computational space can be reduced through selecting only the most important configurations (ways of distributing electrons among molecular orbitals) by a pre-selection algorithm.
 One such algorithm is ADAPT-QSCI [49].
@@
 In this method a simple sampling measurement is performed on an input state prepared by a quantum computer - this is the only step that requires quantum computation.
-The measurement result is used for identifying the most important electron configurations for performing the selected configuration interaction (CI) calculation on classical computers, that is, Hamiltonian diagonalization in the selected $R_{k}$ dimensional subspace $S_{k}=\operatorname{span}\{\ket{r_{1}^{(k)}},\ldots,\ket{r_{R_{k}}^{(k)}}\}$ [40].
+The measurement result is used for identifying the most important electron configurations for performing the selected configuration interaction calculation on classical computers, that is, Hamiltonian diagonalization in the selected $R_{k}$ dimensional subspace $S_{k}=\operatorname{span}\{\ket{r_{1}^{(k)}},\ldots,\ket{r_{R_{k}}^{(k)}}\}$ [49].
 QSCI relies on a quantum computer only for generating the electron configurations via sampling, and the subsequent calculations to output the ground-state energy and the pool operator gradients $h_{j}=\braket{c_{k}}{i[H,P_{j}]}{c_{k}}$ are executed on classical computers333|$\ket{c_{k}}$ is a state corresponding to classical vector $c_{k}=\sum_{l=1}^{R_{k}}(c_{k})_{l}\ket{r_{l}^{(k)}}$, and $i[H,P_{j}]$ is calculated through projecting onto the subspace $S_{k}$ and evaluating the expectation classicaly using the classical vector $c_{k}$.[49]..
@@
 […]
-Figure 1: QCELS circuit [17]
+Figure 1: QCELS circuit [19].
 The QCELS algorithm [19] takes a reference state |$\ket{\psi_{0}}$ and evolves it by the time evolution operator $U(t)=e^{-iHt}$ enclosed within a Hadamard test (as depicted by Figure 1).
@@
 For this challenge, we found that the provided Hamiltonians had an initial Hartree-Fock ground state that retains an overlap with the true ground state - up to $28$ qubits - to extract a good estimate of the energy.
-The time evolution operator was implemented with a first-order Trotter-Suzuki expansion with Hamiltonian truncated to only the Pauli operators with the largest 200 coefficients to reduce gate count.
+The time evolution operator was implemented with a first-order Trotter-Suzuki expansion with Hamiltonian truncated to only the Pauli operators with the largest coefficients to reduce gate count.
 Our procedure for fitting a sum of exponentials to the collected data was to initially fit with a single frequency fitting to
@@
 $f_{\text{fit}}^{(3)}$ $=r_{1}^{(3)}e^{-i\theta_{1}^{(3)}t}+r_{2}^{(3)}e^{-i\theta_{2}^{(3)}t}+(1-r_{1}^{(3)}-r_{2}^{(3)})e^{-i\theta_{3}^{(3)}t}$ (3)
-which were constrained to be smaller in magnitude than the initial fitted frequency $\theta_{1}^{(1)}$ to ensure that they acted as corrections to the initial fit, for the purpose of the challenge we also added a factor $0.8$ to keep the fitting procedure as safe as possible due to the need for it to be automated.
+which were constrained to be smaller in magnitude than the initial fitted frequency $\theta_{1}^{(1)}$ to ensure that they acted as corrections to the initial fit, for the purpose of the challenge we also added a factor to keep the change in $\theta_{1}$ small ensuring that the fitting procedure is as reliable as possible due to the need for it to be automated.
 The size of the amplitude $r_{2}^{(2,3)}$ is also constrained to be smaller than $r_{1}^{(1)}$.
@@
 (c) $f_{\text{fit}}^{(3)}$
-Figure 2: QCELS results from sample data of the orbital rotated Hubbard Hamiltonian on 28 qubits (various parameters)
+Figure 2: QCELS results from sample data of the orbital rotated Hubbard Hamiltonian on 28 qubits (various parameters).
 5 Accelerating Quantum Algorithms with Machine Learning
@@
 Compressing a Hamiltonian in the context of QC reduces the complexity of the Hamiltonian operator but keeps the essential physical properties of the system.
-We compressed by truncating small terms., its exact energy level, and the hyperparameters used to compute this level.
+We compressed by truncating the smallest terms, its exact energy level, and the hyperparameters used to compute this level.
 We use this dataset to train our ML model and predict hyperparameters for new Hamiltonians.
@@
 Moreover, the exact energy level value is irrelevant as we do not use these values directly.
-We are interested in knowing whether certain data (Hamiltonian and some quantum-system-related parameters) lead to further minimised answers (the lowest energy level value).
-This is the relationship between these values.
+We investigate whether specific data (Hamiltonian and quantum-system-related parameters) contribute to further minimising the lowest energy level.
+We are interested in exploring the relationship between these values.
 Therefore, we have used a regressor to predict roughly the lowest energy levels.
-Figure 3: Classical pre-processing phase - training a regressor on smaller systems
+Figure 3: Classical pre-processing phase - training a regressor on smaller systems.
 5.2 AccelerQ
@@
 To ensure the generalisability of AccelerQ and its applicability to various QE algorithms (e.g., ADAPT-QSCI or QCELS), we have developed a versatile wrapper.
-This wrapper encompasses the Hamiltonian, the number of qubits, a parameter indicating whether the computation should be performed classically or on a quantum computer, and the hyperparameters (of both: the QE and ML problems, while we only optimise here those of the QE problem).
+This wrapper encompasses the Hamiltonian, the number of qubits, a parameter indicating whether the computation should be performed classically or on the challenge MPS simulator, and the hyperparameters (of both, the QE and ML problems, while we only optimise here those of the QE problem).
 In our evaluation (see Section 6), we applied AccelerQ to the ADAPT-QSCI and the QCELS algorithms.
 Data Preparation.
-We add open-source commonly used molecular Hamiltonians (H2O, LiH, BeH2, Hemocyanin [2], and H chain) and combine them with the ones provided by the challenge Hamiltonians of $16$ qubits or less, details in [7].
+We use Hamiltonians from [58] and open-source commonly used molecular Hamiltonians (H2O, LiH, BeH2, Hemocyanin [2], and Hydrogen chain) of $16$ qubits.
 The wrapper is run in classical mode while giving a randomly generated set of hyperparameters each time, recording the energy level as the score.
@@
 This stage corresponds to the first of the three steps depicted in Figure 3.
-We utilise a classical eigensolver during data preparation (Subsection 5.1).
+We utilised a quantum eigensolver run on an emulator during data preparation (Subsection 5.1).
 Model Training.
@@
 The rest of the requirements are derived from the rules of the 2024 Quantum Algorithm Grand Challenge (QAGC2024), with one exception: we utilised GPU during the training phase of our models while keeping them relatively small to be able to deploy them on desktop CPU.
-We used the Python library QURI Parts [^5] https://pypi.org/project/quri-parts/ to simulate and access QC.
+We used the Python library QURI Parts [^5] https://pypi.org/project/quri-parts/ to simulate the QC.
 Quantum Algorithm Grand Challenge Description.
@@
 These constraints included a specific machine specification (CPU, 16 GB RAM), a practical time limit of $6\times 10^{5}$ seconds to ensure all solutions could be evaluated within a reasonable timeframe and a shots limit of $10^{7}$ due to the high cost of quantum computer execution.
-Participants were required to make essential use of quantum computers, as the goal was a quantum competition, and current classical algorithms do not scale efficiently for systems with $28$ or more qubits.
+Participants were required to make essential use of quantum computers, as the goal was a quantum competition, and current classical algorithms do not scale efficiently for systems with many ($\gtrsim 50$) qubits.
 Furthermore, participants were instructed to output only the energy level of the ground state, calculated using the quantum algorithm of choice.
-Exact calculations of the ground-state energy using methods like the Bethe ansatz were prohibited, as such methods do not scale well for larger systems ($28$ qubits or more).
+Exact calculations of the ground-state energy using methods like the Bethe ansatz were prohibited, as such methods are either inefficient or do not work for general systems.
 Additionally, participants were not allowed to hard-code initial parameters or hyperparameters to ensure adaptability to new Hamiltonians.
-That constraint ensured that the prosposed solution is generic and not tailored for a specific Hamiltonian or subset of problems.
+That constraint ensures that the proposed solution is generic and not tailored for a specific Hamiltonian or subset of problems.
 To facilitate the automatic evaluation of all participants’ solutions, the contest organisers provided a specific structure for their responses, limiting participants to modifying a single answer file.
@@
 Constraints provided by the contest organisers are particularly relevant for the prediction phase but may not be suitable for the training phase.
-Recognising these limitations, in future iterations of our solution we plan to adopt further a more flexible approach, allowing for greater use of advanced methods and better adaptability, while respecting the original framework and intentions set by the challenge organisers.
+Recognising these limitations, in future iterations of our solution, we plan to adopt a more flexible approach, allowing for greater use of advanced methods and better adaptability while respecting the original framework and intentions set by the challenge organisers.
 6 Evaluation
-We evaluated AccelerQ for its ability to further optimise ADAPT-QSCI and QCELS algorithms by suggesting better hyperparameters tailored per system.
-6.1 Methodolgy
-We demonstrated our idea on the optimisation of hyperparameters of two quantum eigensolver algorithms.
+We evaluated a prototype of AccelerQ for its ability to optimise further ADAPT-QSCI and QCELS algorithms by suggesting better hyperparameters tailored per system, in this report focusing on random search without a crossover operator.
+6.1 Methodology
+We demonstrated our idea by optimising hyperparameters of two quantum eigensolver algorithms.
 We used 4- to 16-qubit systems to train the models.
@@
 We aim to understand:
-RQ1: To what extent can AccelerQ’s optimisation of hyperparameters alone accelerate and improve the efficiency and accuracy of Quantum Eigensolver (QE) algorithms on NISQ devices in terms of runtime, error, and system size?
+RQ1: To what extent can AccelerQ’s optimisation of hyperparameters alone accelerate and improve the efficiency and accuracy of QE algorithms on NISQ devices in terms of error, and system size?
 Given that we extract a model from smaller systems up to 16 qubits, we wish to assess the scalability of these models and to test if we have reduced performance as the system gets larger, that is:
 RQ2: How scalable are machine learning-predicted hyperparameters learned on smaller systems when applied to QE algorithms for Hamiltonian systems with increasing qubit number and complexity?
-In our evaluation, we applied our methodology to 16 larger systems of 20-, 24- and 28-qubit with known lowest energy levels, using the two QE implementations: ADAPT-QSCI and QCELS (Section 4) to answer RQ1 and RQ2.
+In our evaluation, we applied our methodology to 16 larger systems of 20-, 24- and 28-qubit with known lowest energy levels [^6] In the case of the open-source systems, we have only a close estimation using ADAPT-QSCI hyperparameters results as a baseline., using the two QE implementations: ADAPT-QSCI and QCELS (Section 4) to answer RQ1 and RQ2.
 We constructed two models—one for each implementation—using data extracted classically (Section 4) from up to 16-qubit systems.
@@
 The resultant model aimed to predict the optimal hyperparameters for its QE implementation and a system (Hamiltonian).
-The optimisation process took an implementation, its trained model and a system (a problem to solve) as input and returned the optimal hyperparameters for this setup.
+The optimisation process took an implementation, its trained model and a system (a problem to solve) as input, and returned the predicted optimal hyperparameters for this setup.
 Baseline and Parametrisation.
@@
 The hyperparameters were specific per implementation, were optimised, and their performance was compared against the baseline.
-Apart from the hyperparameters, ADAPT-QSCI and QCELS implementations were provided with ham and number_qubits, the input system and the required number of qubits for the corresponding Hamiltonian, and the flag is_classical being set to True during the data collection phase, and either True or False otherwise (classical or QC).
+Apart from the hyperparameters, ADAPT-QSCI and QCELS implementations were provided with ham and number_qubits, the input system and the required number of qubits for the corresponding Hamiltonian, and the flag is_classical being set to True during the data collection phase, and either True or False otherwise (exact classical simulation or MPS).
 We did not optimise this part.
@@
 num_pickup (default: 100) and coeff_cutoff (default: 0.001) parameters control the terms retained or removed from the Hamiltonian compressed representation.
-self_selection indicates whether self-selection is enabled (default: False), which forces the algorithm to work in subspace.
+self_selection indicates whether self-selection is enabled (default: False), which forces the algorithm to work in the subspace with the correct Hamming weight.
 iter_max is the maximum total number of iterations (default: 100).
@@
 final_sampling_shots_coeff is the number of shots for the calculations if the same operator appears twice or the operator parameter is close to zero (default: 5).
-num_precise_gradient is the number of operators from the pool to calculate gradient more precisely (default: 128).
-max_num_converged specifies the maximum number of iterations within atol needed for a solution to be considered converged (default: 2).
+num_precise_gradient is the number of operators from the pool to calculate the gradient with the full Hamiltonian, after an initial approximate calculation is carried out with the truncated Hamiltonian (default: 128).
 reset_ignored_inx_mode specifies after how many iterations previously used operators can be reused (default: 0).
@@
 The default value is 10; n_Z = random.randint(5, 25).
-alpha is a scalar parameter that determines how much smaller the parameters $r_{2}^{(2)}$, $\theta_{2}^{(2)}$ and $r_{2}^{(3)}$, $\theta_{2}^{(3)}$, $\theta_{3}^{(3)}$ are forced to be than (default is $0.8$ ;alpha = random.uniform(0.5, 1)).
+alpha is a scalar parameter that determines how much smaller the parameters introduced in each step of the fitting procedure are than those introduced in the previous step e.g.
+$r_{2}^{(2)}$, $\theta_{2}^{(2)}$ and $r_{2}^{(3)}$, $\theta_{2}^{(3)}$, are bounded by $\alpha r_{1}^{(1)},\pm\alpha\theta_{1}^{(1)}$ (default is $0.8$; alpha = random.uniform(0.5, 1)).
 Using the above, we defined different types of hyperparameter vectors per implementation.
@@
 These results indicate the need for a custom loss function, some feature weighting, and a validation stage with larger datasets, which we leave for future work.
-Table 1: Predicted optimal hyperparameters (A-J col.) with the first row stating the default values for ADAPT-QSCI implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: num_pickup, B: coeff_cutoff, C: self_selection, D: iter_max, E: sampling_shots, F: atol, G: final_sampling_shots_coeff, H: num_precise_gradient, I: max_num_converged, and J: reset_ignored_inx_mode)
-System Size A B C D E F G H I J
-Default - 100 0.001 0 100 100000 1.00E-06 5 128 2 0
-20qubits_00 63636 985 7.64647e-03 0 344658 49458 5.67168e-05 3 77 2 1
-20qubits_01 54717 807 9.41494e-03 1 382109 964853 9.00118e-05 6 79 2 64
-20qubits_02 54723 934 4.56117e-03 0 478268 875460 4.27314e-05 8 294 3 83
-20qubits_03 54710 551 1.13139e-03 1 802530 106374 7.23056e-05 1 161 2 79
-20qubits_04 63636 182 4.06265e-03 0 398085 61950 7.69414e-05 7 194 4 11
-20qubits_05 46 593 8.46933e-04 1 278431 148423 6.66628e-05 7 158 4 58
-24qubits_05 56 67 7.77741e-03 1 655715 676443 3.37411e-05 8 176 2 89
-24qubits_06 56 292 4.00805e-03 0 762863 603105 6.33079e-05 3 159 4 82
-24qubits_07 56 80 4.06535e-03 0 659716 711486 9.45728e-05 4 37 3 0
-24qubits_08 56 215 6.59156e-03 1 885413 291401 1.74522e-05 1 247 4 81
-24qubits_09 56 716 7.29064e-03 0 401350 929361 6.82934e-05 5 257 2 65
-28qubits_01 98700 1000 6.58567e-03 0 615787 300938 1.04770e-05 8 224 4 98
-28qubits_00 98700 197 1.42217e-03 1 460054 209518 7.62522e-05 7 287 3 54
-28qubits_02 98700 531 5.88032e-03 1 402352 258683 4.81220e-06 8 277 2 47
-28qubits_03 98700 629 4.35568e-03 0 429007 82966 3.89304e-05 8 239 4 8
-28qubits_04 98700 698 7.96341e-03 0 221061 633412 7.86532e-05 4 68 4 89
-Table 2: Predicted optimal hyperparameters (A-E cols.) with the first row stating the default values for QCELS implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: delta_t, B: n_Z, C: ham_terms, D: ham_cutoff, and E: alpha)
-System Size A B C D E
-Default - 0.03 10 200 1e-9 0.8
-20qubits_00 63636 1.37929e-01 14 877 6.77275e-03 9.76144e-01
-20qubits_01 54717 2.87833e-01 8 69 9.23310e-04 5.26429e-01
-20qubits_02 54723 1.05366e-01 10 989 4.31098e-04 8.77018e-01
-20qubits_03 54710 1.04235e-01 16 68 6.91101e-03 5.12876e-01
-20qubits_04 63636 1.31645e-01 19 971 2.10235e-03 7.76523e-01
-20qubits_05 46 1.98379e-01 7 187 8.34526e-03 5.33536e-01
-24qubits_05 56 2.08699e-01 24 933 5.75035e-03 6.20185e-01
-24qubits_06 56 1.75486e-01 10 398 7.28888e-03 9.72887e-01
-24qubits_07 56 1.80166e-01 11 286 1.06348e-03 9.75924e-01
-24qubits_08 56 2.72722e-01 22 453 6.67156e-03 8.29813e-01
-24qubits_09 56 2.09069e-01 24 650 3.99377e-03 5.20032e-01
-28qubits_00 98700 7.00174e-03 7 310 6.71154e-03 6.32763e-01
-28qubits_01 98700 1.99713e-01 23 309 8.39957e-03 5.54053e-01
-28qubits_02 98700 5.46760e-02 12 997 1.22706e-03 5.60243e-01
-28qubits_03 98700 4.23353e-02 6 553 8.40543e-03 8.66162e-01
-28qubits_04 98700 2.33005e-01 25 280 1.30071e-04 5.98526e-01
-Table 3: Hamiltonian sizes (Size col.) and the size after compression used when querying the ML Models (Comp.
-col.).
-The column Comp.
-represents the size of the system as utilised in the ML model.
-We marked with an asterisk the open-source molecular Hamiltonians.
-System Size Comp.
-20qubits_00 63636 91
-20qubits_01 54717 287
-20qubits_02 54723 445
-20qubits_03 54710 338
-20qubits_04 63636 65
-20qubits_05(*) 46 46
-System Size Comp.
-24qubits_05(*) 56 48
-24qubits_06(*) 56 46
-24qubits_07(*) 56 50
-24qubits_08(*) 56 51
-24qubits_09(*) 56 47
-System Size Comp.
-28qubits_00 98700 167
-28qubits_01 98700 158
-28qubits_02 98700 98
-28qubits_03 98700 144
-28qubits_04 98700 213
+Table 1: Predicted optimal hyperparameters (A-J col.) with the first row stating the default values for ADAPT-QSCI implementation (A: num_pickup, B: coeff_cutoff, C: self_selection, D: iter_max, E: sampling_shots, F: atol, G: final_sampling_shots_coeff, H: num_precise_gradient, and I: reset_ignored_inx_mode).
+System A B C D E F G H I
+Default 100 0.001 0 100 100000 1.00E-06 5 128 0
+20qubits_00 985 7.64647e-03 0 344658 49458 5.67168e-05 3 77 1
+20qubits_01 807 9.41494e-03 1 382109 964853 9.00118e-05 6 79 64
+20qubits_02 934 4.56117e-03 0 478268 875460 4.27314e-05 8 294 83
+20qubits_03 551 1.13139e-03 1 802530 106374 7.23056e-05 1 161 79
+20qubits_04 182 4.06265e-03 0 398085 61950 7.69414e-05 7 194 11
+20qubits_05 593 8.46933e-04 1 278431 148423 6.66628e-05 7 158 58
+24qubits_05 67 7.77741e-03 1 655715 676443 3.37411e-05 8 176 89
+24qubits_06 292 4.00805e-03 0 762863 603105 6.33079e-05 3 159 82
+24qubits_07 80 4.06535e-03 0 659716 711486 9.45728e-05 4 37 0
+24qubits_08 215 6.59156e-03 1 885413 291401 1.74522e-05 1 247 81
+24qubits_09 716 7.29064e-03 0 401350 929361 6.82934e-05 5 257 65
+28qubits_00 197 1.42217e-03 1 460054 209518 7.62522e-05 7 287 54
+28qubits_01 1000 6.58567e-03 0 615787 300938 1.04770e-05 8 224 98
+28qubits_02 531 5.88032e-03 1 402352 258683 4.81220e-06 8 277 47
+28qubits_03 629 4.35568e-03 0 429007 82966 3.89304e-05 8 239 8
+28qubits_04 698 7.96341e-03 0 221061 633412 7.86532e-05 4 68 89
+Table 2: Predicted optimal hyperparameters (A-E cols.): the approximate size of the Hamiltonian as a vector of terms, A: delta_t, B: n_Z, C: ham_terms, D: ham_cutoff, and E: alpha).
+System A B C D E
+Default 0.03 10 200 1e-9 0.8
+20qubits_00 1.37929e-01 14 877 6.77275e-03 9.76144e-01
+20qubits_01 2.87833e-01 8 69 9.23310e-04 5.26429e-01
+20qubits_02 1.05366e-01 10 989 4.31098e-04 8.77018e-01
+20qubits_03 1.04235e-01 16 68 6.91101e-03 5.12876e-01
+20qubits_04 1.31645e-01 19 971 2.10235e-03 7.76523e-01
+20qubits_05 1.98379e-01 7 187 8.34526e-03 5.33536e-01
+24qubits_05 2.08699e-01 24 933 5.75035e-03 6.20185e-01
+24qubits_06 1.75486e-01 10 398 7.28888e-03 9.72887e-01
+24qubits_07 1.80166e-01 11 286 1.06348e-03 9.75924e-01
+24qubits_08 2.72722e-01 22 453 6.67156e-03 8.29813e-01
+24qubits_09 2.09069e-01 24 650 3.99377e-03 5.20032e-01
+28qubits_00 7.00174e-03 7 310 6.71154e-03 6.32763e-01
+28qubits_01 1.99713e-01 23 309 8.39957e-03 5.54053e-01
+28qubits_02 5.46760e-02 12 997 1.22706e-03 5.60243e-01
+28qubits_03 4.23353e-02 6 553 8.40543e-03 8.66162e-01
+28qubits_04 2.33005e-01 25 280 1.30071e-04 5.98526e-01
+Table 3: Hamiltonian sizes of seven Hamiltonian representations.
+The Hamiltonian sizes in the five rightmost columns are post-Jordan-Wigner Transformation.
+We marked the open-source molecular Hamiltonians with an asterisk.
+(1) FermionOperator term count, in the original Hamiltonian representation (#terms col.); (2) The size of a vector of floating point numbers in NumPy of the compressed, flattened FermionOperator representation for machine learning processing (#floats in ML model col.); (3) initial Hamiltonian term count (Init.
+col.), and its size compressed via cutoff and term selection of (4) default and (5) optimised ADAPT-QSCI, and (6) default and (7) optimised QCELS hyperparameters setup (Round ADAPT-QSCI, Round opt.
+ADAPT-QSCI, Truncated QCELS, and Truncated opt.
+QCELS columns).
+#terms in Hamiltonian (Jordan-Wigner)
+System #terms
+#floats in
+ML model
+Init.
+Round
+ADAPT-QSCI
+Round, Opt.,
+ADAPT-QSCI
+Truncated
+QCELS
+Truncated,
+Opt.,QCELS
+20qubits_00 63636 91 14537 100 985 200 877
+20qubits_01 54717 287 14537 100 807 200 69
+20qubits_02 54723 445 14535 100 934 200 989
+20qubits_03 54710 338 14551 100 551 200 68
+20qubits_04 63636 65 14537 100 182 200 971
+20qubits_05(*) 46 46 67 67 67 66 66
+24qubits_05(*) 56 48 81 81 67 80 80
+24qubits_06(*) 56 46 81 81 81 80 80
+24qubits_07(*) 56 50 81 81 80 80 80
+24qubits_08(*) 56 51 81 81 81 80 80
+24qubits_09(*) 56 47 81 81 81 80 80
+28qubits_00 98700 167 42982 100 197 200 310
+28qubits_01 98700 158 42982 100 1000 200 309
+28qubits_02 98700 98 42982 100 531 200 997
+28qubits_03 98700 144 42982 100 629 200 553
+28qubits_04 98700 213 42982 100 698 200 280
 Optimal Hyperparameters Prediction Results.
 Table 1 and Table 2 show the optimal hyperparameters found by AccelerQ with each Hamiltonian (System col.) employing the two models above.
-Size col.
-is the number of terms in the system.
 The first row shows the default values of the implementation.
 Table 1 shows the ADAPT-QSCI’s optimal hyperparameters, with col.
-A-J represents the predicted optimal hyperparameters values.
-Similarly, Table 2 presents the optimal hyperparameters for the QCELS implementation in col.
+A-I represents the predicted optimal values of the hyperparameters.
+Similarly, Table 2 presents the optimal values for the hyperparameters of the QCELS implementation in col.
 A-E.
@@
 In practice, the iter_max has the same limit as with the default parameters (i.e.
-iter_max=1e7/sampling_shots; col.
+iter_max=1e7/sampling_shots using the values in col.
 E and col.
 D).
-Table 3 shows the size (Size col.) and the compressed size (Comp.
-col.) of the Hamiltonians (System col.) used when querying the ML model during prediction, marking with an asterisk the open-source molecular Hamiltonians.
-Different cutoffs and compression rates were used for ML part than in the QE implementations.
-Consequentially, the sizes recorded in Table 3 are true only for the ML model.
-In Table 3, the number of terms in the open-source molecular Hamiltonians systems ($\leq 46$) is significantly smaller than the challenge Hamiltonians ones ($\geq 54k$), even after compression, the open-source molecular Hamiltonians systems was still much smaller on average.
-This suggests that these systems may differ to some extent.
 The results in Table 1 suggest that the optimisation opted to (sometimes) lower the performance of each iteration, resulting in less precise results per iteration but using more iterations overall by reducing (sometimes) sampling shots and increasing the atol, coeff_cutoff and iter_max.
 However, when the optimisation predicted both iter_max and sampling_shots at maximum values, ignoring their correlation, the execution was capped, resulting in fewer iterations in practice.
-The results in Table 2 show that the optimisation favoured increasing the number of large coefficients (except for 20qubits_01 and 20qubits_03) using larger ham_cutoff values and hence leaving fewer small coefficients.
+The results in Table 2 show that the optimisation favoured increasing the number of large coefficients (except for 20qubits_01 and 20qubits_03) using larger ham_cutoff values and, hence, leaving fewer small coefficients.
 Further, the optimisation selected higher values for n_z and delta_t, with no clear preference for alpha.
-Table 4: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
-System Value Algorithm Error (%) Task Score Aprox. Runtime Itr.
-20qubits_00 -22.046059902 ADAPT-QSCI, Opt. 2.46 -21.503637869562557 15777 s 159
-ADAPT-QSCI, Default 4.52 -21.05032811997207 8211 s 97
-QCELS, Opt. 2.57 -21.48036468528949 14991 s 157
-QCELS, Default 4.47 -21.0598063085247 7936 s 92
-20qubits_01 -22.046059902 ADAPT-QSCI, Opt. 6.26 -20.66624780754983 4476 s 11
-ADAPT-QSCI, Default 4.49 -21.055920933889066 7463 s 84
-QCELS, Opt. 6.30 -20.655947060040507 4327 s 11
-QCELS, Default 4.30 -21.097506543499456 7486 s 82
-20qubits_02 -22.046059902 ADAPT-QSCI, Opt. 5.85 -20.75547153586126 4941 s 12
-ADAPT-QSCI, Default 4.55 -21.042520869027147 6370 s 65
-QCELS, Opt. 5.85 -20.756855659992507 4789 s 12
-QCELS, Default 4.62 -21.02776101198928 6152 s 65
-20qubits_03 -22.046059902 ADAPT-QSCI, Opt. 3.37 -21.302881787984727 8631 s 95
-ADAPT-QSCI, Default 4.63 -21.024636461355875 6682 s 77
-QCELS, Opt. 3.27 -21.32597699721156 8575 s 95
-QCELS, Default 4.59 -21.03455240312227 6293 s 66
-20qubits_04 -22.046059902 ADAPT-QSCI, Opt. 3.31 -21.316521673460805 11004 s 121
-ADAPT-QSCI, Default 4.64 -21.024150729181112 7595 s 88
-QCELS, Opt. 3.34 -21.308808719093236 9807 s 108
-QCELS, Default 4.57 -21.03798918026242 7039 s 83
-28qubits_00 -30.748822808 ADAPT-QSCI, Opt. 6.35 -28.7974854444735 10812 s 48
-ADAPT-QSCI, Default 6.42 -28.77456507714568 10345 s 61
-QCELS, Opt. 6.35 -28.794794630409676 11076 s 48
-QCELS, Default 6.40 -28.779806038153325 10989 s 65
-28qubits_01 -30.748822808 ADAPT-QSCI, Opt. 6.39 -28.783766574436143 10024 s 34
-ADAPT-QSCI, Default 6.47 -28.758306955348417 11789 s 77
-QCELS, Opt. 6.41 -28.777776815200454 10096 s 34
-QCELS, Default 6.46 -28.763094962187317 11249 s 74
-28qubits_02 -30.748822808 ADAPT-QSCI, Opt. 6.41 -28.778922785865355 11373 s 39
-ADAPT-QSCI, Default 6.68 -28.69374047078425 9479 s 53
-QCELS, Opt. 6.39 -28.7835710030780 10666 s 39
-QCELS, Default 6.66 -28.70067301821086 9509 s 57
-28qubits_03 -30.748822808 ADAPT-QSCI, Opt. 6.25 -28.827497968717516 11248 s 51
-ADAPT-QSCI, Default 6.47 -28.75988277903682 12548 s 85
-QCELS, Opt. 6.18 -28.848400001238723 11644 s 53
-QCELS, Default 6.66 -28.699830076240456 11168 s 70
-28qubits_04 -30.748822808 ADAPT-QSCI, Opt. 6.79 -28.662288396044573 6757 s 16
-ADAPT-QSCI, Default 6.77 -28.667008949938953 8603 s 43
-QCELS, Opt. 6.78 -28.66351868143365 7463 s 16
-QCELS, Default 6.70 -28.689855023621764 5909 s 45
-Table 5: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
-System Algorithm Task Score Aprox. Runtime Itr.
-20qubits05 ADAPT-QSCI, Opt. -13.120566186315502 3568 s 27
-ADAPT-QSCI, Default -14.000000000000009 686 s 9
-QCELS, Opt. -12.968777862300158 3719 s 26
-QCELS, Default -14.000000000000004 681 s 9
-24qubits05 ADAPT-QSCI, Opt. -9.034894851637292 3029 s 11
-ADAPT-QSCI, Default -10.99319108738972 7053 s 35
-QCELS, Opt. -8.622105677825234 4741 s 8
-QCELS, Default -9.039194172208862 5367 s 35
-24qubits06 ADAPT-QSCI, Opt. -9.460283005059706 4718 s 8
-ADAPT-QSCI, Default -9.460283005059702 644 s 6
-QCELS, Opt. -9.460283005059704 4467 s 8
-QCELS, Default -9.4602830050597 663 s 6
-24qubits07 ADAPT-QSCI, Opt. -7.34876884915856 4637 s 12
-ADAPT-QSCI, Default -10.400021274590904 8117 s 37
-QCELS, Opt. -7.932213481832702 3977 s 10
-QCELS, Default -9.973622581884937 5117 s 36
-24qubits08 ADAPT-QSCI, Opt. -6.274613702752347 4560 s 35
-ADAPT-QSCI, Default -6.034877031453911 1516 s 17
-QCELS, Opt. -5.598926935684615 4779 s 35
-QCELS, Default -6.034877031453906 1585 s 17
-24qubits09 ADAPT-QSCI, Opt. -13.136457290516125 4660 s 7
-ADAPT-QSCI, Default -13.136457290516129 649 s 7
-QCELS, Opt. -13.13645729051612 3941 s 6
-QCELS, Default -13.13645729051613 686 s 7
+The hyperparameters affect the number of terms considered during the QE implementation executions, leading to different cutoffs, the number of terms to select and compression rates used for the ML model and the QE implementation.
+Table 3 presents an analysis of these sizes under different choices of values of the hyperparameters for the ML model and both QE implementations.
+The Hamiltonian sizes in the five rightmost columns in Table 3 are post-Jordan-Wigner Transformation.
+Per Hamiltonian (System col.), we recorded the FermionOperator term count in the original Hamiltonian representation (#terms col.), the initial Hamiltonian term count in Jordan Wigner form (Init.
+col.), and this representation’s size when compressed via cutoff and term selection of the default and optimised ADAPT-QSCI, and the default and optimised QCELS hyperparameters setup (Round ADAPT-QSCI col., Round, opt., ADAPT-QSCI col., Truncated QCELS col.
+and Truncated, opt., QCELS col., respectively).
+The column #floats in ML model is the size of a vector of floating point numbers in NumPy of the compressed, flattened FermionOperator representation for machine learning processing.
+We construct this vector as follows.
+The Hamiltonian, initially expressed as a collection of terms with associated coefficients, is transformed into a flat numerical vector to streamline processing and analysis.
+This transformation involves extracting the coefficients, filtering out those with negligible contributions, and selecting key elements significant enough to influence the overall Hamiltonian.
+The resulting coefficients are then organised into a consistent and compact format.
+This representation retains the essential numerical properties of the Hamiltonian while ensuring uniformity and suitability for downstream processing.
+Lastly, we marked the open-source molecular Hamiltonians with an asterisk in Table 3.
+The source of the Hamiltonians is described in Subsection 6.2.
+In Table 3, the number of terms in the open-source molecular Hamiltonian systems ($\leq 46$) is smaller by several orders of magnitude compared to the challenge Hamiltonians ($\geq 54k$).
+Even after compression, the open-source molecular Hamiltonians systems were still much smaller on average.
+This suggests that these systems may differ to some extent.
+The next section examines the differences in the percentage of error achieved by each system, both with and without optimised hyperparameters, and in relation to their sizes.
 Results of Execution with Different Hyperparameters.
 Table 4 and Table 5 summarise the results of executing ADAPT-QSCI and QCELS on a quantum simulator with the default and optimal hyperparameters in Table 1 and Table 2.
-We evaluated predictions on several Hamiltonians of 20-, 24-, and 28-qubit systems (System col.) with a known solution (Value col.).
+We evaluated predictions on several Hamiltonians of 20-, 24- and 28-qubit systems (System col.); we excluded the results of QCELS on 28 qubits from Table 4 and discuss these in detail in the discussion paragraph below.
 Table 4 presents the result obtained with the challenge Hamiltonians, while Table 5 is the results using the open-source molecular Hamiltonians.
-Both tables present, for each algorithm and set of hyperparameters (Algorithm col.), the task score (Task Score col.), the approximate run time (Approx.
-Runtime col.), and the number of iterations completed (Itr.
+Both tables present, for each algorithm and set of hyperparameters (Algorithm col.), the task score (Task Score col.), and the number of iterations completed (Itr.
 col.).
 For Table 4, the true results (Value col.) are known [58], and we computed the relative error (Error (%) col.) per experiment.
-We ran each experiment several times and selected the best result.
+Note that in total, we evaluated our prototype framework using 16 systems, each tested with two implementations (ADAPT-QSCI and QCELS) under default and optimised hyperparameters, resulting in 16×2×2=64 experiments.
+Each experiment was repeated 10 times, and the best (lowest) result from these repetitions was selected [^9] For QCELS, the implementation frequently crashed in the machine during our experiments, likely due to insufficient resources, making it challenging to obtain results.
+To mitigate this, we adopted a resilient execution strategy, running additional repetitions and collecting results only from the first 10 non-crashing executions.
+In some cases, we were fortunate, and all executions completed without crashing; in these cases, we excluded the last machine’s results to maintain consistency across experiments..
+While the time per repeat is excluded from this paper and tables due to unreliable recording on our host machine, timing logs are available in our Zenodo record [9].
+Table 4: The challenge Hamiltonians: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed when the result was obtained.).
+System Value Algorithm Error (%) Task Score Itr.
+20qubits_00 -22.046059902 ADAPT-QSCI, Opt. 2.42 -21.5134481079129 147
+ADAPT-QSCI, Default 4.28 -21.102120951277936 85
+QCELS, Opt. 8.17 -20.244706524965135 13
+QCELS, Default 10.2 -19.79974257492709 9
+20qubits_01 -22.046059902 ADAPT-QSCI, Opt. 6.27 -20.66347471739218 10
+ADAPT-QSCI, Default 4.65 -21.02122671542096 78
+QCELS, Opt. 0.83 -21.86260739785393 7
+QCELS, Default 7.64 -20.361475620508653 9
+20qubits_02 -22.046059902 ADAPT-QSCI, Opt. 5.79 -20.76850360872923 11
+ADAPT-QSCI, Default 4.28 -21.101500354754492 65
+QCELS, Opt. 8.43 -20.18686453079191 9
+QCELS, Default 8.47 -20.178731347808355 9
+20qubits_03 -22.046059902 ADAPT-QSCI, Opt. 3.24 -21.33253115024456 93
+ADAPT-QSCI, Default 4.49 -21.055181221011736 79
+QCELS, Opt. 6.79 -20.548184887148853 15
+QCELS, Default 8.81 -20.104190198375786 9
+20qubits_04 -22.046059902 ADAPT-QSCI, Opt. 3.30 -21.31780361442556 126
+ADAPT-QSCI, Default 4.59 -21.034981116395308 83
+QCELS, Opt. 8.53 -20.16538489337923 15
+QCELS, Default 7.05 -20.49223412640405 9
+28qubits_00 -30.748822808 ADAPT-QSCI, Opt. 6.33 -28.80238180353666 45
+ADAPT-QSCI, Default 6.44 -28.76881293227412 58
+28qubits_01 -30.748822808 ADAPT-QSCI, Opt. 6.37 -28.788828203287586 33
+ADAPT-QSCI, Default 6.36 -28.79342085633283 76
+28qubits_02 -30.748822808 ADAPT-QSCI, Opt. 6.38 -28.788448328499076 38
+ADAPT-QSCI, Default 6.60 -28.720685906958145 59
+28qubits_03 -30.748822808 ADAPT-QSCI, Opt. 6.11 -28.871013405346638 56
+ADAPT-QSCI, Default 6.44 -28.767725361229388 75
+28qubits_04 -30.748822808 ADAPT-QSCI, Opt. 6.78 -28.66362404354463 15
+ADAPT-QSCI, Default 6.71 -28.68471298123107 37
+Table 5: Open-source molecular Hamiltonians: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr.
+is Iterations Completed).
+We are not providing the exact ground state energy values for reference due to the significant computational difficulty in obtaining them for these systems.
+Instead, we present the energy values obtained from ADAPT-QSCI, which serve as a reliable lower bound for the ground state energy.
+For QCELS, we fixed the task score by a constant factor that was missing in the original computation.
+System Algorithm Task Score Itr.
+20qubits05 ADAPT-QSCI, Opt. -13.308108432842927 26
+ADAPT-QSCI, Default -14.000000000000021 9
+QCELS, Opt. -13.602690623110092 6
+QCELS, Default -13.57495914514915 9
+24qubits05 ADAPT-QSCI, Opt. -8.965984985383866 10
+ADAPT-QSCI, Default -9.667517470048752 40
+QCELS, Opt. -7.987619347414383 23
+QCELS, Default -8.23727545736091 9
+24qubits06 ADAPT-QSCI, Opt. -9.460283005059704 7
+ADAPT-QSCI, Default -9.460283005059702 5
+QCELS, Opt. -9.35586373343566 9
+QCELS, Default -9.342675357790028 9
+24qubits07 ADAPT-QSCI, Opt. -7.570508389218553 12
+ADAPT-QSCI, Default -9.965785682971944 27
+QCELS, Opt. -7.7996489422244103 10
+QCELS, Default -7.6227840132533573 9
+24qubits08 ADAPT-QSCI, Opt. -6.277553305948219 27
+ADAPT-QSCI, Default -7.74244754268563 30
+QCELS, Opt. -5.1411838474119161 21
+QCELS, Default -5.2184341011492506 9
+24qubits09 ADAPT-QSCI, Opt. -13.136457290516129 5
+ADAPT-QSCI, Default -13.390021054767033 14
+QCELS, Opt. -13.13645729051612 23
+QCELS, Default -12.799039886601325 9
 In Table 4, the optimised hyperparameters tended to achieve better results than the default.
-The optimised ADAPT-QSCI found the best solution for 4 systems, optimised QCELS for 3, default QCELS for 2, and default ADAPT-QSCI for 1 out of 10 tested systems.
-For 20-qubit systems, optimised ADAPT-QSCI had the lowest error at 4.25%, followed by optimised QCELS at 4.27%.
-Default QCELS and ADAPT-QSCI had average error rates of 4.51% and 4.57%, respectively.
-For 28-qubit systems, optimised QCELS had the lowest error at 6.42%, followed by optimised ADAPT-QSCI at 6.44%, default ADAPT-QSCI at 6.56%, and default QCELS at 6.58%.
+The optimised ADAPT-QSCI found the best solution for 6 systems, optimised QCELS for 1, default ADAPT-QSCI for 3, and default QCELS for none, out of 10 tested systems.
+For 20-qubit systems, optimised ADAPT-QSCI had the lowest error at 4.2%, followed by default ADAPT-QSCI at 4.46%.
+Optimised and default QCELS had average error rates of 6.55% and 8.43%, respectively.
+For 28-qubit systems, optimised ADAPT-QSCI had the lowest error at 6.39%, followed by default ADAPT-QSCI at 6.51%.
 Table 5 shows that default parameters performed better.
-The default ADAPT-QSCI found the best solution for 3 systems, followed by optimised ADAPT-QSCI for 2, and default QCELS for 1 out of 6 tested systems.
-Answer to RQ1.
+The default ADAPT-QSCI found the best solution for 5 systems, followed by optimised ADAPT-QSCI for 1, out of 6 tested systems.
+RQ1 Answer.
 The results indicate a limited ability to improve solely through hyperparameter optimisation, though the optimised hyperparameters performed better in Table 4.
-A more refined model based on Hamiltonian characteristics is required to properly assess the optimisation of hyperparameters’ impact on accelerating and improving QE efficiency and accuracy.
-In detail: (20-qubit) Both implementations, with default or optimised hyperparameters, had similar iteration counts with the default setup consuming less time on average.
-However, optimised setups performed worse when the number of iterations was too low.
-(24-qubit) Time consumption increased for both ADAPT-QSCI and QCELS with optimised hyperparameters and used fewer iterations when the error percentage was higher than in the default setup.
-(28-qubit) The number of iterations was halved on average, with optimised setups requiring 38 compared to 62-64 iterations with default setups.
-For all systems, optimised hyperparameters used 42-44 iterations versus 53-56 for the default setups, with some increase in time consumption for the optimised setup.
-9 systems performed better with optimised setups, while 7 performed better with the default setups, mostly from Table 5.
-Answer to RQ2.
-Our understanding of AccelerQ’s scalability remains somewhat limited: 80% of the 28-qubit systems’ scores improved with optimised setup, while only 45% of the 20- and 24-qubit systems (5 out of 11) showed improvement.
-This suggests other factors, like Hamiltonian nature, hyperparameter types, padding, or even data proportions (per source of Hamiltonians), affect scalability and require further investigation.
-In detail: Table 4 shows that optimised hyperparameters performed better for the challenge Hamiltonians.
-We observed the opposite for the open-source molecular Hamiltonians Table 5, particularly for optimised QCELS.
-The challenge Hamiltonian systems contributed 8.4% of ADAPT-QSCI’s and 37.9% of QCELS’s training data.
-There is also a correlation between the models: when QCELS with optimised hyperparameters outperformed the default setup, ADAPT-QSCI typically did as well, and vice versa.
-This suggests that the data each type of Hamiltonian contributed might not be the primary factor.
-Lastly, the open-source molecular Hamiltonians systems had fewer terms used in the ML phase in comparison to the challenge Hamiltonians, which may affect the model’s ability to generalise [^8] Open-source molecular Hamiltonians averaged 48.2 elements, compared to 200.6 for the challenge Hamiltonians, Table 3..
+A more refined model based on Hamiltonian characteristics is required to assess the optimisation of hyperparameters’ impact on accelerating and improving QE efficiency and accuracy.
+RQ1 answer in detail: (20-qubit) Optimised setups showed larger error variability (0.83%–8.53%) than default setups (4.28%–10.2%).
+While the QCELS optimised version achieved the lowest error (0.83%) with just seven iterations, the ADAPT-QSCI optimised version performed better in other cases.
+Both implementations, with default vs optimised hyperparameters, had similar iteration counts, with QCELS requiring fewer iterations in general.
+However, optimised setups mostly performed worse when the number of iterations was too low.
+(24-qubit) ADAPT-QSCI default outperformed the rest of the configurations, with no particular further observations.
+It suggests that optimisation did not enhance accuracy for systems whose Hamiltonians have a relatively small number of terms.
+(28-qubit) The number of iterations was almost halved on average, with optimised setups requiring 37 compared to 61 iterations with default setups, suggesting potential acceleration benefits.
+Furthermore, optimised ADAPT-QSCI, when outperforming the default setup, achieved lower errors—often by an order of magnitude—while default setups only had marginal advantages (0.x vs 0.0x error differences).
+This suggests that hyperparameter optimisation can yield substantial accuracy gains despite not always outperforming default settings.
+For all systems, ADAPT-QSCI optimised hyperparameters used 41 iterations versus 51 for the default setup, with the opposite trend for QCELS, whose runs with optimised hyperparameters used 14 iterations versus 9 with the default setup.
+Eight systems performed better with optimised setups, mostly from Table 4, while eight performed better with the default setups, mostly from Table 5.
+RQ2 Answer.
+Our understanding of AccelerQ’s scalability remains somewhat limited: 70% of the challenge Hamiltonians’ scores improved with optimised setup along with a general decrease of the error in general but showed little improvement in open-source molecular Hamiltonians.
+This suggests that other factors, like Hamiltonian nature, hyperparameter types, padding, or even data proportions (per source of Hamiltonians), affect scalability and require further investigation.
+RQ2 answer in detail: We considered here two factors: (1) training data size and (2) prior hyperparameter optimisation.
+The challenge Hamiltonian systems contributed 8.4% of ADAPT-QSCI’s and 37.9% of QCELS’s training data, with QCELS using much shorter vectors.
+Open-source molecular Hamiltonians averaged 48 elements, compared to 200.6 for the challenge Hamiltonian (Table 3, #floats in ML model col.).
+Table 4 shows that optimised hyperparameters performed better for the challenge Hamiltonians.
+We observed the opposite for the open-source molecular Hamiltonians in Table 5.
+This suggests that training data quantity alone is not the primary factor but rather a feature dimension—as open-source molecular Hamiltonians had fewer terms in the ML phase, which may affect the model’s ability to generalise.
+Finally, the default QCELS performed the worst, as its original hyperparameters had not been previously optimised by other methods, possibly explaining its weaker performance in comparison to the challenge algorithm, ADAPT-QSCI.
+Discussion.
+Analysis of the scalability of the optimisation procedure is hampered by the limitations of classical simulators.
+For larger system sizes, the time to evaluate a large circuit can quickly become prohibitive for running many iterations of algorithms, even with access to large compute resources and sufficient memory to store the system’s state.
+These challenges prevented the successful evaluation of the QCELS algorithm at $28$ qubits with a strong damping effect on the oscillations of time evolving expectation value.
+Possible causes of such an effect could include the Trotter error [68, 66] introduced by the implementation of the Hamiltonian evolution unitary or the truncation of entanglement between qubits in the Matrix Product State simulator [79, 4, 78, 50].
+In the previous version of the pre-print the results provided for QCELS at $28$ qubits were based on data that was strongly damped and so was not well fit by our function of three complex exponentials with a dominant frequency of the ground state energy and instead a high frequency fit with a period of approximately the distance between data points was found, giving the appearance of a reasonable result.
+In figure 4(a), the decay of the oscillation can be observed in the data as well as the erroneous fit.
+In figure 4(b), we forbid a fit with a period smaller than the spacing of the data, however, the decay of the oscillations continues to prevent a good fit to the data.
+In comparison, the adapt-QSCI algorithm continues to work well up to $28$ qubits, with its formulation as a classical solver acting in a subspace defined from the measurements on a quantum computer being well suited to working at moderate system sizes.
+It also works well at mitigating the impact of inexact quantum evolution from noise on a quantum device or error in the classical simulation of the quantum algorithm.
+(a) The original high-frequency fitting procedure with a period of approximately the spacing between the data.
+The damping of the oscillations is observed in the data collected from the simulated quantum algorithm.
+(b) An updated fitting procedure which prevents high frequency fits, but does not significantly improve the results in this case because of the error in the collected time evolution data.
 7 Conclusion
@@
 Acknowledgments.
-We thank CloudLab [18] for the platform and infrastructure support that enabled the experiment resulting in the construction of two models for accelerating quantum algorithms.
+Authors are listed in alphabetical order.
+We thank CloudLab [22] for the platform and infrastructure support that enabled the experiment, resulting in the construction of two models for accelerating quantum algorithms.
 Sophie Fortz, Connor Lenihan and Avner Bensoussan are partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1.
-Sophie Fortz and Avner Bensoussan are partially funded by the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009.
-Sophie is partially funded by Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1 and by the QAssure project from Innovate UK.
-Also King’s Quantum grants provided by King’s College London are gratefully acknowledged.
+Sophie Fortz and Avner Bensoussan are partially funded by the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009 - Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1.
+Sophie Fortz is partially funded by the QAssure project from Innovate UK.
+Also, King’s Quantum grants provided by King’s College London are gratefully acknowledged.